### v1

# Gradient Boosting Reinforcement Learning

Abstract
Neural networks (NN) achieve remarkable results in various tasks, but lack key characteristics: interpretability, support for categorical features, and lightweight implementations suitable for edge devices. While ongoing efforts aim to address these challenges, Gradient Boosting Trees (GBT) inherently meet these requirements. As a result, GBTs have become the go-to method for supervised learning tasks in many real-world applications and competitions. However, their application in online learning scenarios, notably in reinforcement learning (RL), has been limited. In this work, we bridge this gap by introducing Gradient-Boosting RL (GBRL), a framework that extends the advantages of GBT to the RL domain. Using the GBRL framework, we implement various actor-critic algorithms and compare their performance with their NN counterparts. Inspired by shared backbones in NN we introduce a tree-sharing approach for policy and value functions with distinct learning rates, enhancing learning efficiency over millions of interactions. GBRL achieves competitive performance across a diverse array of tasks, excelling in domains with structured or categorical features. Additionally, we present a high-performance, GPU-accelerated implementation that integrates seamlessly with widely-used RL libraries (available at https://github.com/NVlabs/gbrl). GBRL expands the toolkit for RL practitioners, demonstrating the viability and promise of GBT within the RL paradigm, particularly in domains characterized by structured or categorical features.
1 Introduction
Reinforcement Learning (RL) has shown great promise in various domains that involve sequential decision making. However, many real-world tasks, such as inventory management, traffic signal optimization, network optimization, resource allocation, and robotics, are represented by structured observations with categorical or mixed data types. These tasks can benefit significantly from deployment and training on edge devices due to resource constraints. Moreover, interpretability is crucial in these applications for regulatory reasons and for trust in the decision-making process. Current neural network (NN) based solutions struggle with interpretability, handling categorical data, and supporting light implementations suitable for low-compute devices.
Gradient Boosting Trees (GBT) is a powerful ensemble method extensively used in supervised learning due to its simplicity, accuracy, interpretability, and natural handling of structured and categorical data. Frameworks such as XGBoost [7], LightGBM [20], and CatBoost [36] have become integral in applications spanning finance [49], healthcare [54, 27, 43], and competitive data science [6]. Despite their successes, GBT has seen limited application in RL. This is primarily because traditional GBT libraries are designed for static datasets with predefined labels, contrasting with the dynamic nature of RL. The distribution shift in both input (state) and output (reward) poses significant challenges for the direct application of GBT in RL. Moreover, there is a notable lack of benchmarks or environments tailored for structured data, further hindering progress in this area.
In this paper, we introduce Gradient Boosting Reinforcement Learning (GBRL), a GBT framework tailored for RL. Our contributions are:
  1. 1.
    GBT for RL. We demonstrate the viability and potential of GBT as function approximators in RL. We present GBT-based implementations of PPO, A2C, and AWR, and show that GBRL is competitive with NNs across a range of environments. In addition, similarly to supervised learning, GBRL outperforms NNs on categorical tasks (see Figure 1).
  2. 2.
    Tree-based Actor-Critic architecture. Inspired by shared architectures in NN-based actor-critic (AC), we introduce a GBT-based AC architecture. This reduces the memory and computational requirements by sharing a common ensemble structure for both the policy and value. This approach significantly reduces runtime compared to existing GBT frameworks, thus removing the barrier to solving complex, high-dimensional RL tasks with millions of interactions.
  3. 3.
    Modern GBT-based RL library  [^1] The GBRL core library is available at https://github.com/NVlabs/gbrl. The actor-critic implementations based on GBRL are available at https://github.com/NVlabs/gbrl_sb3.. We provide a CUDA-based [33] hardware-accelerated GBT framework optimized for RL. GBRL is designed to work as part of a broader system and seamlessly integrates with popular repositories such as Stable-baselines3 [39]. This new tool offers practitioners a powerful option for exploring GBT in RL settings.
Figure 1: PPO GBRL vs PPO NN. Aggregated mean and standard deviation of the normalized average reward for the final 100 episodes. Rewards were normalized as: $\text{reward}_{\text{norm}}=\frac{\text{reward}}{\text{reward}_{\text{max}\{\text{NN},\text{GBRL}\}}}$ per environment and then aggregated across each domain.
2 Related Work
Gradient boosted trees.
Recent advances have extended GBT’s capabilities beyond traditional regression and classification. In ranking problems, GBT has been used to directly optimize ranking metrics, as demonstrated by frameworks like StochasticRank [51] and recent advancements explored in Lyzhin et al. [26]. Additionally, GBT offer probabilistic predictions through frameworks like NGBoost [11], enabling uncertainty quantification [28]. The connection between GBT and Gaussian Processes [52, 45] offers further possibilities for uncertainty-aware modeling. Recently, Ivanov and Prokhorenkova [18] modeled graph-structured data by combining GBT with graph neural networks.
Despite their versatility, applying GBT in RL remains a relatively less explored area. Several works have employed GBT as a function approximator within off-policy RL methods, including its use in Q-learning [1] and in bandit settings to learn inverse propensity scores [24]. Recently, Brukhim et al. [5] proposed a boosting framework for RL where a base policy class is incrementally enhanced using linear combinations and nonlinear transformations. However, these previous works have not yet demonstrated the scalability and effectiveness in complex, high-dimensional RL environments requiring extensive interactions. In this work, we show how to adapt the framework of GBT to successfully solve large-scale RL problems.
Interpretability.
Due to the inherent non-linearities, NNs are challenging to interpret and require sophisticated methods to do so. Interpreting NNs often involves either approximation with simpler models such as decision trees or using gradient-based techniques, which require additional forward and backward passes [16, 9, 44, 3, 37]. On the other hand, interpretability methods for GBT can take advantage of the structure of a decision tree for high speed, efficiency, and accuracy [25, 10].
Structured and categorical data.
Previous work in RL has predominantly focused on using NNs due to their ability to capture complex patterns in high-dimensional data. Techniques such as Q-learning and AC methods have advanced significantly, demonstrating success in tasks involving raw sensory inputs like images, text, and audio. However, NNs that perform well on structured and categorical data typically have very specialized architectures [19, 46, 14, 2] and are not standard multi-layer perceptrons (MLPs) that are often used in many RL tasks and algorithms [34]. Even with these specialized architectures, Gradient Boosting Trees (GBT) often perform equally or better on structured and categorical datasets [19, 31, 14, 15].
Policy optimization through functional gradient ascent.
In this approach, the policy is parameterized by a growing linear combination of functions [29]. Each linear addition represents the functional gradient with respect to current parameters. Kersting and Driessens [21] demonstrated the direct optimization of policies using the policy gradient theorem [48]. Similarly, Scherrer and Geist [41] proposed a functional gradient ascent approach as a local policy search algorithm. While these works lay theoretical groundwork, practical results on complex, high-dimensional RL environments have not been shown. To adapt GBT’s to RL, we leverage the framework of functional gradient ascent. This combination enables a seamless integration of GBRL directly into existing RL optimization packages, such as Stable-baselines [39].
3 Preliminaries
We begin by introducing Markov Decision Processes (MDPs) and the AC schema. Then, we introduce GBT. In the following section, we show how to combine both of these paradigms into GBRL.
3.1 Markov Decision Process
We consider a fully observable infinite-horizon Markov decision process (MDP) characterized by the tuple $(\mathcal{S},\mathcal{A},P,\mathcal{R})$. At each step, the agent observes a state $\operatorname{\mathbf{s}}\in\mathcal{S}$ and samples an action $\operatorname{\mathbf{a}}\in\mathcal{A}$ from its policy $\pi(\operatorname{\mathbf{s}},\operatorname{\mathbf{a}})$. Performing the action causes the system to transition to a new state $\operatorname{\mathbf{s}}^{\prime}$ based on the transition probabilities $P(\operatorname{\mathbf{s}}^{\prime}|\operatorname{\mathbf{s}},\operatorname{\mathbf{a}})$, and the agent receives a reward $\operatorname{\mathbf{r}}\sim\mathcal{R}(\operatorname{\mathbf{s}},\operatorname{\mathbf{a}})$. The objective is to find an optimal policy $\pi^{*}$ that maximizes the expected discounted reward $J(\pi)=\mathbb{E}[\sum_{t=0}^{\infty}\gamma^{t}\operatorname{\mathbf{r}}_{t}]$, with a discount factor $\gamma\in[0,1)$.
The action-value function $Q_{\pi}(\operatorname{\mathbf{s}},\operatorname{\mathbf{a}}):=\mathbb{E}_{\pi}[\sum_{t^{\prime}=0}^{\infty}\gamma^{t^{\prime}}\mathcal{R}(\operatorname{\mathbf{s}}_{t+t^{\prime}},\operatorname{\mathbf{a}}_{t+t^{\prime}})|\operatorname{\mathbf{s}}_{t}=\operatorname{\mathbf{s}},\operatorname{\mathbf{a}}_{t}=\operatorname{\mathbf{a}}]$ estimates the expected returns of performing action $\operatorname{\mathbf{a}}$ in state $\operatorname{\mathbf{s}}$ and then acting according to $\pi$. Additionally, the value function $V_{\pi}(\operatorname{\mathbf{s}}):=\mathbb{E}_{\pi}[\sum_{t^{\prime}=0}^{\infty}\gamma^{t^{\prime}}\mathcal{R}(\operatorname{\mathbf{s}}_{t+t^{\prime}},\operatorname{\mathbf{a}}_{t+t^{\prime}})|\operatorname{\mathbf{s}}_{t}=\operatorname{\mathbf{s}}]$, predicts the expected return starting from state $\operatorname{\mathbf{s}}$ and acting according to $\pi$. Finally, the advantage function $A_{\pi}(\operatorname{\mathbf{s}},\operatorname{\mathbf{a}}):=Q_{\pi}(\operatorname{\mathbf{s}},\operatorname{\mathbf{a}})-V_{\pi}(\operatorname{\mathbf{s}})$, indicates the expected relative benefit of performing action $\operatorname{\mathbf{a}}$ over acting according to $\pi$.
3.2 Actor-Critic Reinforcement Learning
Actor-critic methods are a common method to solve the objective $J(\pi)$. They learn both the policy and value. In the GBRL framework, we extend three common AC algorithms to support GBT-based function approximators.
A2C [32] is a synchronous, on-policy AC algorithm designed to improve learning stability. The critic learns a value function, $V(\operatorname{\mathbf{s}})$, used to estimate the advantage. This advantage is incorporated into the policy gradient updates, reducing variance and leading to smoother learning. The policy is updated using the following gradient: $\nabla_{\theta}J(\pi_{\theta})=\mathbb{E}[\nabla_{\theta}\log\pi_{\theta}(\operatorname{\mathbf{a}}|\operatorname{\mathbf{s}})A(\operatorname{\mathbf{s}},\operatorname{\mathbf{a}})]$.
PPO [42] extends A2C by improving stability. This is achieved through constrained policy update steps using a clipped surrogate objective. This prevents drastic policy changes and leads to smoother learning. To achieve this, PPO solves the following objective: $\nabla_{\theta}J(\pi_{\theta})=\mathbb{E}[\nabla_{\theta}\text{clip}(\frac{\log\pi_{\theta}(\operatorname{\mathbf{a}}|\operatorname{\mathbf{s}})}{\log\pi_{\theta_{\text{old}}}(\operatorname{\mathbf{a}}|\operatorname{\mathbf{s}})},1-\epsilon,1+\epsilon)A(\operatorname{\mathbf{s}},\operatorname{\mathbf{a}})]$. Additionally, PPO enhances sample efficiency by performing multiple optimization steps on each collected rollout.
AWR [35] is an off-policy AC algorithm. Provided a dataset $\mathcal{D}$, AWR updates both the policy and the value through supervised learning. This dataset can be pre-defined and fixed (offline), or continually updated using the agents experience (replay buffer). At each training iteration $k$, AWR solves the following two regression problems:
$V_{k}$ $=\operatorname*{arg\,min}_{V}\mathbb{E}_{\operatorname{\mathbf{s}},\operatorname{\mathbf{a}}\sim\mathcal{D}}[\|G(\operatorname{\mathbf{s}},\operatorname{\mathbf{a}})-V(\operatorname{\mathbf{s}})\|_{2}^{2}]\,,\,\pi_{k+1}=\operatorname*{arg\,max}_{\pi}\mathbb{E}_{\operatorname{\mathbf{s}},\operatorname{\mathbf{a}}\sim\mathcal{D}}[\log\pi(\operatorname{\mathbf{a}}|\operatorname{\mathbf{s}})\exp(\frac{1}{\beta}A_{k}(\operatorname{\mathbf{s}},\operatorname{\mathbf{a}}))]\,,$
where $G(\operatorname{\mathbf{s}},\operatorname{\mathbf{a}})$ represents the monte-carlo estimate or TD($\lambda$) of the expected return [47].
3.3 Gradient Boosting Trees as Functional Gradient Descent
Gradient boosting trees (GBT) [12] are a non-parametric machine learning technique that combines decision tree ensembles with functional gradient descent [30]. GBT iteratively minimizes the expected loss $L(F(\mathbf{x}))=\mathbb{E}_{\mathbf{x},\mathbf{y}}[L(\mathbf{y},F(\mathbf{x}))]$ over a dataset $D=\{(\mathbf{x}_{i},\mathbf{y}_{i})\}_{i=1}^{N}$. A GBT model, $F_{K}$, predicts outputs using $K$ additive trees as follows:
$F_{K}(\mathbf{x}_{i})=F_{0}+\sum_{k=1}^{K}\epsilon h_{k}(\mathbf{x}_{i})\,,$ (1)
where $\epsilon$ is the learning rate, $F_{0}$ is the base learner, and each $h_{k}$ is an independent regression tree partitioning the feature space.
In the context of functional gradient descent, the objective is to minimize the expected loss $L(F(\mathbf{x}))=\mathbb{E}_{\mathbf{x},\mathbf{y}}[L(\mathbf{y},F(\mathbf{x}))]$ with respect to the functional $F$. Here, a functional $F:\mathcal{H}\rightarrow\mathbb{R}$ maps a function space to real numbers. A GBT model can be viewed as a functional $F$ that maps a linear combination of binary decision trees to outputs: $F:\text{lin}(\mathcal{H})\rightarrow\mathbb{R}^{D}$, where $\mathcal{H}$ is the decision tree function class.
We start with an initial model, $F_{0}$, and iteratively add trees to $F$ to minimize the expected loss. Similar to parametric gradient descent, at each iteration $k$, we minimize the loss by taking a step in the direction of the functional gradient $g^{i}_{k}:=\nabla_{F_{k-1}}L(\mathbf{y}_{i},F_{k-1}(\mathbf{x}_{i}))$. However, we are constrained to gradient directions within $\mathcal{H}$. Thus, we project the gradient $g_{k}$ into a decision tree by solving:
$h_{k}=\arg\min_{h}\|-\epsilon g_{k}-h(\mathbf{x})\|_{2}^{2}\,.$ (2)
4 Gradient Boosting Reinforcement Learning
Figure 2: The GBRL framework. The actor’s policy and critic’s value function are parameterized by $\theta_{k}$. For example, $\theta_{k}=[\mu(\operatorname{\mathbf{s}}),\sigma(\operatorname{\mathbf{s}}),V(\operatorname{\mathbf{s}})]$ for a Gaussian policy. $\theta_{k}$ is calculated by summing all the outputs of trees in the ensemble. Starting from $\theta_{0}$, at each training iteration, GBRL collects a rollout and computes the gradient $\nabla_{\theta_{0}}J$. This gradient is then used to fit the next tree added to the ensemble, which is updated to $\theta_{1}$. This process repeats with each iteration fitting a new tree, refining the parameterization, and expanding the ensemble towards $\theta_{k}\approx\theta_{0}+\epsilon\sum^{k-1}_{m=0}\nabla_{\theta_{m}}J$, an approximated scaled sum of gradients with respect to past parametrizations.
In this work, we extend the framework of GBT to support AC algorithms in the task of RL. The objective in RL is to optimize the return $J$, the cumulative reward an agent receives. Unlike in supervised learning, the target predictions are unknown a priori. RL agents learn through trial and error. Good actions are reinforced by taking a step in the direction of the gradient $\nabla_{\pi}J$. This formulation aligns perfectly with functional gradient ascent; thus, in GBRL, we optimize the objective directly over the decision tree function class. This is achieved by iteratively growing the ensemble of trees $\{h_{i}\}$. The ensemble outputs $\theta$, representing AC parameters such as the policy $\pi$ and the value function. For example, $\theta=[\mu(\operatorname{\mathbf{s}}),\sigma(\operatorname{\mathbf{s}}),V(\operatorname{\mathbf{s}})]$ for a Gaussian policy. At each iteration, a new tree $h_{k}$, constructed to minimize the distance to $\nabla_{\theta_{k-1}}J$, is added to the ensemble. Here, The resulting method is an application of GBT as a functional gradient optimizer $\theta_{k}\approx\theta_{0}+\epsilon\sum^{k-1}_{m=0}\nabla_{\theta_{m}}J$.
However, RL presents unique challenges for GBT. RL involves a nonstationary state distribution and inherent online learning, causing gradients to vary in magnitude and direction. Large gradients in unfavorable directions risk destabilizing training or leading to catastrophic forgetting. Moreover, feedback in RL is provided through interactions with the environment and is not available a priori. This contrasts with supervised learning settings, where gradients decrease with boosting iterations, and targets are predefined. As a result, many of the key features that traditional GBT libraries rely on are not suitable. For example, GOSS [20], categorical feature encoding [36], early-stopping signals, pruning methods [53], and strategies to tackle online learning [55].
To address these challenges, we employ appropriate tools from the NN and GBT literature, such as batch learning [13, 40] to update the ensemble. At each boosting iteration, we fit a decision tree on a random batch sampled with replacement from the experience buffer. This approach helps handle non-stationary distributions and improve stability by focusing on different parts of the state space, allowing beneficial gradient directions to accumulate and minimizing the impact of detrimental ones. Additionally, GBRL fits gradients directly to optimize objectives, whereas traditional GBT methods require targets and need workarounds to utilize gradients effectively.
A common theme in AC algorithms is to utilize a shared approximation for the actor and the critic. We adopt this approach in GBRL, constructing trees where each leaf provides two predictions. GBRL predicts both the policy (distribution over actions) and the value estimate. The internal structure of the tree is shared, providing a single feature representation for both objectives and significantly reducing memory and computational bottlenecks. Accordingly, in GBRL we apply differentiated learning rates to the policy and value outputs during prediction, effectively optimizing distinct objectives within this shared structure. We present the full algorithm in Algorithm 1 and diagram in Figure 2.
Algorithm 1 Gradient Boosting for Reinforcement Learning (GBRL)
1:Initialize: $\theta_{0}$, $\epsilon_{\text{actor}}$, $\epsilon_{\text{critic}}$, experience buffer $\mathcal{B}$, total training iterations $K$, number of updates $U$, batch size $N$, $k\leftarrow 1$
2:while $k<K$ do
3:     Collect trajectory $\tau^{(k)}=(\operatorname{\mathbf{s}}_{0},\operatorname{\mathbf{a}}_{0},\ldots,\operatorname{\mathbf{s}}_{T},\operatorname{\mathbf{a}}_{T})^{(k)}$ and rewards $(\operatorname{\mathbf{r}}_{0},\ldots,\operatorname{\mathbf{r}}_{T})^{(k)}$ using $\pi_{\theta_{k-1}}$
4:     Add trajectory $\tau^{(k)}$ and rewards to the experience buffer $\mathcal{B}$
5:     for each update $u=1$ to $U$ do
6:         Sample a batch from the experience buffer $\mathcal{B}$
7:         Compute gradients $g$ according to AC algorithm (e.g., PPO, A2C, AWR)
8:         Construct dataset $D=\{(\operatorname{\mathbf{s}}_{n},g_{n})\}_{n=0}^{N}$ and fit a decision tree $h_{k}$
9:         for each dimension $d=0$ to $D$ do
10:              if $0\leq d<D$ then
11:                  Update $\theta^{(d)}_{k}=\theta^{(d)}_{k-1}+\epsilon_{\text{actor}}h_{k}(\operatorname{\mathbf{s}})$
12:              else
13:                  Update $\theta^{(d)}_{k}=\theta^{(d)}_{k-1}+\epsilon_{\text{critic}}h_{k}(\operatorname{\mathbf{s}})$
14:         $k\leftarrow k+1$
15:Output: AC parameters $\theta^{(d)}_{K}(\operatorname{\mathbf{s}})$ for $d=0,1,\ldots,D$
5 Experiments
Our experiments aim to answer the following questions:
  1. 1.
    GBT as RL Function Approximator: Can GBT-based AC algorithms effectively solve complex high-dimensional RL tasks?
  2. 2.
    Comparison to NNs: How does GBRL compare with NN-based training in various RL algorithms?
  3. 3.
    Benefits in Categorical Domains: Do the benefits of GBT in supervised learning transfer to the realm of RL?
  4. 4.
    Comparison to Traditional GBT libraries: Can we use traditional GBT libraries instead of GBRL for RL tasks?
  5. 5.
    Evaluating the shared AC architecture: How does sharing the tree structure between the actor and the critic impact performance?
We implemented GBT-based versions of A2C, PPO, and AWR within Stable Baselines3. We refer to our implementations as PPO GBRL, A2C GBRL, and AWR GBRL. We evaluated GBRL against the equivalent NN implementations. Where available, we utilize hyperparameters from RL Baselines3 Zoo [38]; otherwise, we optimize the hyperparameters for specific environments. The AWR NN implementation is based on the original paper [35].
We conducted experiments on a range of RL domains. We test classic control tasks, high-dimensional vectorized problems, and finally categorical tasks. We use 5 random seeds per experiment on a single NVIDIA V100-32GB GPU. We present the cummulative non-discounted reward, averaged across the last 100 episodes. We normalize the plots for simple visual comparison between GBRL and the corresponding NN implementations. The normalized score is computed as $\text{score}_{\text{norm}}=\frac{\text{score}_{\text{GBRL}}-\text{score}_{\text{NN}}}{\text{score}_{\text{max}\{\text{NN},\text{GBRL}\}}-\text{score}_{\text{min}\{\text{NN},\text{GBRL}\}}}$. We provide the full learning curves, implementation details, compute resources, un-normalized numerical results, and hyperparameters in the supplementary material.
Classic Enviroments.
We evaluate GBRL’s ability to solve classic RL tasks using Continuous-Control and Box2D environments, provided via Gym [50]. We trained agents for 1M steps (1.5M for LunarLander-v2) and provide the results in Figure 3. For exact values, refer to Table 2.
Considering the algorithmic objective, we observe that GBRL and NN present similar performance when optimized using PPO. In contrast, the other methods demonstrate inconclusive results. In certain environments, such as MountainCar, GBRL outperforms NN with all AC methods. On the other hand, in Pendulum NN is better.
Figure 3: Continuous-Control and Box2D environments. Normalized comparison between GBRL and NN. PPO, the best performing method, shows similar performance with GBRL and NN function classes.
High-Dimensional Vectorized Environments.
The decision-tree function class operates on individual features at each step. Consequently, this function class is not well-suited for handling pixel-based representations, which require more complex feature interactions. Therefore, we evaluted GBRL in the Football [22] and Atari RAM [4] domains. These offer high-dimensional vectorized representations. We trained agents in both environments for 10 million timesteps. The complete results are reported in Tables 3 and 4 and illustrated in Figure 4.
The results portray the following phenomenon. While both tasks may seem similar, there is a distinct difference. The features in the football domain are manually constructed and represent identifiable information, such as the location of the ball and the players. However, the Atari RAM domain provides a flattened view of the system RAM, which is unstructured.
At their core, binary decision trees are if-else clauses. This function class is naturally suited to work with structured data. These insights are emphasized in the football domain. Here, PPO GBRL greatly outperforms PPO NN across most environments and exhibits equivalent performance on the rest. In addition, as Atari RAM is unstructured, we observe that, as can be expected, in most cases GBRL underperforms NN, except for AWR. However, AWR NN underperformed considerably compared to the other NN implementations.
(a) Football Academy environments.
(b) Atari-ramNoFrameskip-v4 environments.
Figure 4: High-Dimensional Vectorized Environments. GBRL outperforms NN on the structured Football domain using PPO. NN outperforms on unstructured tasks, such as Atari RAM.
Categorical Environments.
The football experiment suggests that GBRL outperforms when assigned structured data. Here, we focus on categorical environments. This is a regime where GBT excels in supervised learning. In these experiments, we evaluated the MiniGrid domain [8]. It consists of 2D grid worlds with goal-oriented tasks that require object interaction. We trained in PutNear, FourRooms, and Fetch tasks for 10M timesteps, matching the reported PPO NN in RL Baselines3 Zoo. We trained the remaining environments for 1M timesteps. We give the results in Figure 5. For exact numbers, see Table 5.
Figure 5: MiniGrid environments. GBRL combined with the PPO backend outperforms NN across a range of categorical environments.
In MiniGrid, GBRL outperforms or is on-par with NN in most tasks. Specifically, PPO GBRL is significantly better than PPO NN. We observe the same trend when comparing between environments. These results emphasize that GBRL is a strong candidate for problems characterized by structured data, specifically when using PPO as the algorithmic backend.
GBRL vs Traditional GBT Libraries.
Here, we compare GBRL with Catboost and XGBoost. We focus on the PPO variant. When comparing to the standard libraries, we utilize their built-in options for incremental learning, vectorized leaves, and custom loss functions. As both CatBoost and XGBoost do not support differential learning rates, we used separate ensembles for the actor and the critic. For the comparison, we use the CartPole-v1 environment, training for 1M steps. The results are shown in Figure 6.
As seen, standard GBT libraries are unable to solve RL tasks in a realistic timeframe. GBRL, however, efficiently solves the task while also remaining competitive with NN across a range of environments.
Figure 6: Comparing to standard GBT libraries. CatBoost and XGBoost are intractable in RL. Specifically, CatBoost lacks GPU support for custom losses, leading to low FPS and early termination.
Evaluating the shared AC architecture.
Finally, we evaluated the benefits of using a shared AC architecture by training PPO GBRL on three MiniGrid environments. We train agents with shared and non-shared architectures for 10M timesteps and compare the score, GPU memory usage, and FPS. We provide the aggregated results in Figure 7, and environment-specific breakdowns in the supplementary.
The benefit of the shared structure is clear both in terms of GPU memory consumption and computation speed. By sharing the tree structure, GBRL requires less than half the memory and almost triples the training FPS. This is achieved without any negative performance on the resulting policy, as seen in the reward plot.
Figure 7: Shared Actor-Critic. Sharing the tree structure significantly increases training efficiency and memory, without impacting on the score.
Result summary.
The performance of GBRL varied across RL algorithms, but environments like MiniGrid highlight the potential advantages of using GBT in RL. The results suggest that GBT’s strengths in handling structured and categorical data from supervised learning can effectively transfer to the RL domain. Conversely, GBRL underperformed in Atari-RAM environments, indicating that certain environments, characterized by unstructured observations, are less suited for GBTs.
The results can be explained by the findings of Grinsztajn et al. [15], which suggest that NNs have an inductive bias toward overly smooth solutions and that MLP-like architectures are not robust to uninformative features. The optimal solutions for Atari-RAM might be smoother, which could explain the better performance of NNs. On the other hand, McElfresh et al. [31] argue that GBT outperforms NNs on ‘irregular’ datasets. Tree-based models excel in handling irregular patterns and categorical data, aligning with GBRL’s success in environments like MiniGrid.
Comparing different algorithmic backbones, we find PPO to be the strongest. PPO GBRL excelled in the MiniGrid and Football domains, and performed comparably with NN in classic control tasks. PPO GBRL’s success can be attributed to its alignment with GBRL’s incremental learning strategy. On the other hand, A2C’s single gradient update per rollout may limit its effectiveness and contribute to its underwhelming performance in many environments. Similarly, AWR’s design for multiple sample updates results in very large ensembles, creating a trade-off between large, slow, and memory-intensive ensembles, and lighter, less performant versions.
6 Conclusion
Historically, RL practitioners have relied on tabular, linear, and NN-based function approximators. But, GBT, a widely successful tool in supervised learning, has been absent from this toolbox. We present a method for effectively integrating it into RL and demonstrate domains where it excels compared to NNs. GBRL is a step toward solutions that are more interpretable, well suited for real-world tasks with structured data, or capable of deployment on low-compute devices.
The choice of an RL method depends on the task characteristics: tabular and linear approaches are suitable for small state spaces or simple mappings, while NNs handle complex relationships in unstructured data. GBT thrives in complex, yet structured environments. In such cases, we observe the advantage of GBRL over NNs, reflecting its already known benefits in supervised learning.
A crucial component of GBRL is our efficient adaptation of GBT for AC methods, which allows the simultaneous optimization of distinct objectives. We optimized this approach for large-scale ensembles using GPU acceleration (CUDA). Furthermore, GBRL integrates seamlessly with existing RL libraries, promoting ease of use and adoption.
7 Limitations and Future Directions
In this work, we integrated the highly popular GBT, typically used in supervised learning, into RL. Our results show that GBT is competitive across a range of problems. However, we identified several limitations and compelling areas for further research. First, a significant challenge lies in the continuous generation of trees. As the policy improves through numerous updates, the size of the ensemble increases. This unbounded growth has implications for memory usage, computational efficiency, and the feasibility of online real-time adaptation. The problem is exacerbated by off-policy methods that build many trees per sample. Moreover, the redundancy of trees, especially those from early stages, suggests that the final policy could be represented with a much smaller ensemble. Consequently, developing strategies for tree pruning, ensemble compression, or dynamically managing ensemble size could offer crucial optimizations without compromising performance.
Another key challenge lies in effectively integrating GBT with additional state-of-the-art RL algorithms such as DDPG [23] or SAC [17]. These require differentiable Q-functions to update the policy. Since GBTs are not differentiable, new solutions are needed to incorporate them into these algorithms. One such possible direction can be probabilistic trees, where each node represents the probability of traversing the graph.
Appendix
This appendix provides supplementary materials that support the findings and methodologies discussed in the main text. It is organized into four sections to present the full experiment results, implementation details, hyperparameters used during the experiments, training progression plots, and experimental plots, respectively. These materials offer detailed insights into the research process and outcomes, facilitating a deeper understanding and replication of the study.
Appendix A Implementaion Details and Hyperparameters
Included in this section are implementation details, information regarding compute resources, and tables containing the hyperparameters used in our experiments enabling the reproducibility of our results. Table 1 lists GBRL hyperparameters for all experiments.
A.1 Environments
The Football domain consists of a vectorized 115-dimensional observation space that summarizes the main aspects of the game and 19 discrete actions. We focus on its academy scenarios, which present situational tasks involving scoring a single goal. A standard reward of +1 is granted for scoring, and we employed the "Checkpoints" shaped reward structure. This structure provides additional points as the agent moves closer towards the goal, with a maximum reward of 2 per scenario. The Atari-ram environment consists of a vectorized 128-dimensional observational space representing the 128 byte RAM state and up to 18 discrete actions. We trained agents in both domains for 10M timesteps.
The MiniGrid environment [8] is a 2D grid world with goal-oriented tasks requiring object interaction. The observation space consists of a 7x7 image representing the grid, a mission string, and the agent’s direction. Each tile in the observed image contains a 3D tuple dictating an object’s color, type, and state. All MiniGrid tasks emit a reward of +1 for successful completion and 0 otherwise.
We trained our NN-based agents on a flattened observation space using the built-in one-hot wrapper. For GBRL agents, we generated a 51-dimensional categorical observational space by encoding each unique tile tuple as a categorical string to represent the observed image. Categorical features were added for the agent’s direction (up, left, right, down) and missions. All agents were trained for 1M timesteps, except for PutNear, FourRooms, and Fetch tasks, which were trained for 10M based on the reported values for PPO NN in RL Baselines3 Zoo.
batch size clip range ent coef gae lambda gamma
num epochs
num steps
num envs
policy lr
value lr
Acrobot 512 0.2 0.0 0.94 0.99 20 128 16 0.16 0.034
CartPole 64 0.2 0.0 0.8 0.98 1 128 8 0.029 0.015
LunarLander 256 0.2 0.0033 0.98 0.999 20 512 16 0.031 0.003
MountainCar 256 0.2 0.033 0.98 0.999 20 512 16 0.031 0.003
MountainCar Continuous
256 0.2 0.033 0.98 0.999 20 512 16 0.031 0.003
Pendulum 512 0.2 0.0 0.93 0.91 20 256 16 0.031 0.013
Football 512 0.2 0.0 0.95 0.998 10 256 16 0.033 0.006
Atari-Ram 64 0.92 8e-5 0.95 0.99 4 512 16 0.05 0.002
MiniGrid 512 0.2 0.0 0.95 0.99 20 256 16 0.17 0.01
(a) PPO. For continuous action spaces we used log std init = -2 and log std lr = lin_0.0017. We utilized gradient norm clipping for Gym environments. Specifically, 10 for the value gradients and 150 for the policy gradients.
ent coef gae lambda gamma
num steps
num envs
policy lr
value lr
log std init
log std lr
Acrobot 0.0 1 0.99 8 4 0.79 0.031 - -
CartPole 0.0 1 0.99 8 16 0.13 0.047 - -
LunarLander 0.0 1 0.995 5 32 0.16 0.04 - -
MountainCar 0.0 1 0.99 8 16 0.64 0.032 - -
MountainCar Continuous
0.0 1 0.995 128 16 0.0008 2.8e-6 0 0.0004
Pendulum 0.0 0.9 0.9 10 32 0.003 0.056 -2 0.00018
Football 0.0004 0.95 0.998 128 8 0.87 0.017 - -
Atari-Ram 0.0009 0.95 0.993 128 8 0.17 0.013 - -
MiniGrid 0.0 0.95 0.99 10 128 0.34 0.039 - -
(b) A2C
batch size
ent coef gae lambda gamma
train freq
gradient steps
num envs
policy lr
value lr
log std init
log std lr
Acrobot 1024 0.0 0.95 0.99 2000 150 1 0.05 0.1 - -
CartPole 1024 0.0 0.95 0.99 2000 150 1 0.05 0.1 - -
LunarLander 1024 0.0 0.95 0.99 2000 150 1 0.05 0.1 - -
MountainCar 64 0.0 0.95 0.99 2000 150 1 0.64 0.032 - -
MountainCar Continuous
64 0.0 0.95 0.99 2000 150 1 0.089 0.083 -2 lin_0.0017
Pendulum 1024 0.0 0.9 0.9 1000 50 1 0.003 0.07 -2 0.0005
Football 512 0.03 0.95 0.99 750 10 1 0.09 0.00048 - -
Atari-Ram 1024 0.0 0.95 0.993 2000 50 1 0.0779 0.0048 - -
MiniGrid 1024 0.0 0.95 0.99 1500 25/100* 1 0.0075 0.005 - -
(c) AWR. For all envs, buffer size = 50,000, $\beta$ = 0.05. *MiniGrid environments used 100 gradient steps for tasks trained for 1M steps, and 25 gradient steps for tasks trained for 10M steps, for a reduced tree size.
Table 1: GBRL hyperparameters - NN represented by an MLP with two hidden layers.
A.2 Compute Resources
All experiments were done on the NVIDIA NGC platform on a single NVIDIA V100-32GB GPU per experiment. Training time and compute requirements vary between algorithms and according to hyperparameters. The number of boosting iterations has the largest impact on both runtime and memory. GBRL experimental runs required from 1GB to 24GB of GPU memory. Moreover, runtime varied from 20 minutes for 1M timesteps training on classic environments and up to 5 days for 10M timesteps on Atari-ram. NN experimental runs required up to 3GB of GPU memory and runtime ranged from 10 minutes and up to 3 days. The total compute time for all experiments combined was approximately 1800 GPU hours. Additionally, the research project involved preliminary experiments and hyperparameter tuning, which required an estimated additional 168 GPU hours.
Appendix B Detailed Results Tables
This section contains tables presenting the mean and standard deviation of the average episode reward for the final 100 episodes within each experiment. More specifically, Table 2 presents results for Continuous Control & Block2D environments, Tables 3 and 4 present results for the high-dimensional vectorized environments, and Table 5 presents results for the categorical environments.
Table 2: Continuous-Control and Box2D environments: Average episode reward for the final 100 episodes.
Acrobot CartPole LunarLander MountainCar MountainCar Continuous Pendulum-v1
NN: A2C $-82.27\pm 3.29$ $500.00\pm 0.0$ $-43.01\pm 106.26$ $-148.90\pm 24.10$ $92.66\pm 0.32$ $-183.64\pm 22.32$
GBRL: A2C $-90.73\pm 2.98$ $500.00\pm 0.0$ $47.93\pm 41.00$ $-124.42\pm 5.74$ $93.15\pm 1.19$ $-538.83\pm 66.25$
NN: AWR $-102.53\pm 57.25$ $500.00\pm 0.0$ $282.48\pm 1.96$ $-160.65\pm 53.97$ $18.93\pm 42.34$ $-159.64\pm 9.42$
GBRL: AWR $-118.12\pm 33.54$ $497.54\pm 3.11$ $76.03\pm 56.62$ $-146.68\pm 24.53$ $44.38\pm 45.94$ $-1257.61\pm 98.10$
NN: PPO $-74.83\pm 1.22$ $500.00\pm 0.0$ $261.73\pm 6.93$ $-115.53\pm 1.39$ $85.81\pm 7.51$ $-249.31\pm 60.00$
GBRL: PPO $-87.82\pm 2.16$ $500.00\pm 0.0$ $248.72\pm 59.10$ $-110.55\pm 15.60$ $89.42\pm 5.73$ $-246.89\pm 20.61$
Table 3: Football Academy environments: Average episode reward for the final 100 episodes.
3 vs 1 with keeper Corner Counterattack Easy Counterattack Hard Empty Goal Empty Goal Close
NN: A2C $1.78\pm 0.10$ $1.00\pm 0.17$ $1.58\pm 0.35$ $1.43\pm 0.17$ $1.93\pm 0.05$ $2.0\pm 0.0$
GBRL: A2C $1.59\pm 0.17$ $1.01\pm 0.07$ $1.11\pm 0.14$ $1.00\pm 0.05$ $1.81\pm 0.03$ $2.00\pm 0.00$
NN: AWR $1.50\pm 0.37$ $1.01\pm 0.04$ $1.59\pm 0.36$ $1.18\pm 0.21$ $1.90\pm 0.08$ $1.92\pm 0.17$
GBRL: AWR $1.66\pm 0.34$ $0.92\pm 0.05$ $0.95\pm 0.05$ $0.92\pm 0.05$ $1.93\pm 0.07$ $2.0\pm 0.0$
NN: PPO $1.61\pm 0.05$ $0.95\pm 0.02$ $1.43\pm 0.15$ $1.23\pm 0.18$ $1.98\pm 0.01$ $1.99\pm 0.00$
GBRL: PPO $1.63\pm 0.19$ $1.05\pm 0.20$ $1.64\pm 0.09$ $1.23\pm 0.07$ $1.84\pm 0.06$ $2.0\pm 0.0$
Pass & Shoot keeper Run Pass & Shoot keeper Run to Score Run to score w/ keeper Single Goal vs Lazy
NN: A2C $1.41\pm 0.37$ $1.77\pm 0.08$ $1.87\pm 0.12$ $1.25\pm 0.23$ $1.65\pm 0.04$
GBRL: A2C $1.60\pm 0.21$ $1.60\pm 0.14$ $1.82\pm 0.10$ $1.15\pm 0.08$ $1.31\pm 0.11$
NN: AWR $1.26\pm 0.46$ $1.15\pm 0.14$ $1.81\pm 0.14$ $1.25\pm 0.34$ $1.28\pm 0.27$
GBRL: AWR $1.35\pm 0.37$ $1.53\pm 0.40$ $1.98\pm 0.01$ $0.99\pm 0.16$ $1.03\pm 0.12$
NN: PPO $1.31\pm 0.13$ $1.64\pm 0.16$ $1.91\pm 0.09$ $1.13\pm 0.06$ $1.68\pm 0.09$
GBRL: PPO $1.87\pm 0.09$ $1.85\pm 0.08$ $1.83\pm 0.04$ $1.95\pm 0.02$ $1.73\pm 0.06$
Table 4: Atari-ramNoFrameskip-v4 environments: Average episode reward for the final 100 episodes.
Alien Amidar Asteroids Breakout Gopher
NN: A2C $1802.24\pm 323.12$ $304.62\pm 55.61$ $2770.46\pm 271.97$ $76.69\pm 30.08$ $3533.84\pm 118.50$
GBRL: A2C $595.08\pm 43.51$ $48.71\pm 14.65$ $1402.66\pm 161.67$ $11.52\pm 2.34$ $502.20\pm 341.88$
NN: AWR $739.82\pm 303.06$ $86.32\pm 40.16$ $2308.68\pm 257.72$ $26.57\pm 9.91$ $1471.93\pm 716.65$
GBRL: AWR $829.99\pm 166.48$ $125.53\pm 25.25$ $1592.63\pm 109.96$ $17.32\pm 1.89$ $913.06\pm 79.95$
NN: PPO $1555.32\pm 107.59$ $310.93\pm 80.13$ $2309.46\pm 145.66$ $32.88\pm 15.74$ $2507.84\pm 108.37$
GBRL: PPO $1163.86\pm 76.54$ $186.32\pm 50.63$ $1514.34\pm 317.46$ $19.96\pm 1.93$ $1215.04\pm 81.01$
Kangaroo Krull MsPacman Pong SpaceInvaders
NN: A2C $2137.6\pm 425.64$ $9325.38\pm 777.12$ $2007.64\pm 116.52$ $15.39\pm 4.26$ $462.30\pm 35.56$
GBRL: A2C $948.8\pm 483.80$ $5291.4\pm 433.35$ $989.68\pm 100.02$ $-12.80\pm 11.10$ $265.36\pm 44.64$
NN: AWR $1214.8\pm 313.42$ $4519.78\pm 522.11$ $892.31\pm 289.36$ $-10.25\pm 2.11$ $842.00\pm 130.51$
GBRL: AWR $1809.26\pm 37.51$ $6419.26\pm 387.76$ $1641.84\pm 284.19$ $-11.68\pm 3.79$ $397.85\pm 566.38$
NN: PPO $2487.4\pm 829.65$ $9167.3\pm 294.30$ $2069.22\pm 202.48$ $18.50\pm 1.60$ $479.77\pm 65.07$
GBRL: PPO $2160.8\pm 826.92$ $6888.66\pm 756.18$ $2069.22\pm 538.62$ $15.40\pm 6.55$ $434.84\pm 31.83$
Table 5: MiniGrid environments: Average episode reward for the final 100 episodes.
DoorKey-5x5 Empty-Random-5x5 Fetch-5x5-N2 FourRooms GoToDoor-5x5
NN: A2C $0.96\pm 0.00$ $0.77\pm 0.42$ $0.43\pm 0.03$ $0.62\pm 0.19$ $0.05\pm 0.04$
GBRL: A2C $0.96\pm 0.00$ $0.96\pm 0.00$ $0.62\pm 0.02$ $0.51\pm 0.07$ $0.78\pm 0.02$
NN: AWR $0.57\pm 0.52$ $0.96\pm 0.00$ $0.90\pm 0.26$ $0.19\pm 0.12$ $0.95\pm 0.01$
GBRL: AWR $0.96\pm 0.00$ $0.97\pm 0.00$ $0.95\pm 0.01$ $0.54\pm 0.05$ $0.94\pm 0.01$
NN: PPO $0.78\pm 0.40$ $0.96\pm 0.00$ $0.89\pm 0.03$ $0.53\pm 0.03$ $0.60\pm 0.06$
GBRL: PPO $0.96\pm 0.00$ $0.96\pm 0.00$ $0.96\pm 0.01$ $0.56\pm 0.04$ $0.96\pm 0.00$
KeyCorridorS3R1 PutNear-6x6-N2 RedBlueDoors-6x6 Unlock
NN: A2C $0.75\pm 0.42$ $0.01\pm 0.00$ $0.30\pm 0.22$ $0.77\pm 0.43$
GBRL: A2C $0.39\pm 0.48$ $0.18\pm 0.018$ $0.0\pm 0$ 0.0 $0.90\pm 0.09$
NN: AWR $0.93\pm 0.00$ $0.60\pm 0.13$ $0.83\pm 0.00$ $0.96\pm 0.00$
GBRL: AWR $0.94\pm 0.00$ $0.36\pm 0.01$ $0.84\pm 0.03$ $0.95\pm 0.00$
NN: PPO $0.76\pm 0.42$ $0.001\pm 0.00$ $0.17\pm 0.40$ $0.97\pm 0.00$
GBRL: PPO $0.95\pm 0.00$ $0.44\pm 0.19$ $0.88\pm 0.02$ $0.97\pm 0.00$
Appendix C Training Plots
This section presents learning curves depicting model performance throughout the training phase. Figures 8, 9, 10 and 11 show the training reward as a function of environment steps of the agents trained in the experiments. The column order is: A2C, AWR, and PPO.
Figure 8: Classic Control and Box2D environments: Training reward as a function of environment steps.
Figure 9: Football Academy environments: Training reward as a function of environment step.
Figure 10: Atari-ramNoFrameskip-v4 environments: Training reward as a function of environment step.
Figure 11: MiniGrid environments: Training reward as a function of environment step.
Figure 12: Sharing actor critic tree structure significantly increases efficiency while retraining similar performance. Training reward, GPU memory usage, and FPS, are compared across 10M environment (5 seeds, 3 MiniGrid environments)

### v2

@@
 Abstract
-Neural networks (NN) achieve remarkable results in various tasks, but lack key characteristics: interpretability, support for categorical features, and lightweight implementations suitable for edge devices.
-While ongoing efforts aim to address these challenges, Gradient Boosting Trees (GBT) inherently meet these requirements.
-As a result, GBTs have become the go-to method for supervised learning tasks in many real-world applications and competitions.
-However, their application in online learning scenarios, notably in reinforcement learning (RL), has been limited.
-In this work, we bridge this gap by introducing Gradient-Boosting RL (GBRL), a framework that extends the advantages of GBT to the RL domain.
-Using the GBRL framework, we implement various actor-critic algorithms and compare their performance with their NN counterparts.
-Inspired by shared backbones in NN we introduce a tree-sharing approach for policy and value functions with distinct learning rates, enhancing learning efficiency over millions of interactions.
-GBRL achieves competitive performance across a diverse array of tasks, excelling in domains with structured or categorical features.
-Additionally, we present a high-performance, GPU-accelerated implementation that integrates seamlessly with widely-used RL libraries (available at https://github.com/NVlabs/gbrl).
-GBRL expands the toolkit for RL practitioners, demonstrating the viability and promise of GBT within the RL paradigm, particularly in domains characterized by structured or categorical features.
+We present Gradient Boosting Reinforcement Learning (GBRL), a framework that adapts the strengths of gradient boosting trees (GBT) to reinforcement learning (RL) tasks.
+While neural networks (NNs) have become the de facto choice for RL, they face significant challenges with structured and categorical features and tend to generalize poorly to out-of-distribution samples.
+These are challenges for which GBTs have traditionally excelled in supervised learning.
+However, GBT’s application in RL has been limited.
+The design of traditional GBT libraries is optimized for static datasets with fixed labels, making them incompatible with RL’s dynamic nature, where both state distributions and reward signals evolve during training.
+GBRL overcomes this limitation by continuously interleaving tree construction with environment interaction.
+Through extensive experiments, we demonstrate that GBRL outperforms NNs in domains with structured observations and categorical features while maintaining competitive performance on standard continuous control benchmarks.
+Like its supervised learning counterpart, GBRL demonstrates superior robustness to out-of-distribution samples and better handles irregular state-action relationships.
+Machine Learning, Reinforcement Learning, Gradient Boosting Trees, Gradient Boosting, ICML
 1 Introduction
-Reinforcement Learning (RL) has shown great promise in various domains that involve sequential decision making.
-However, many real-world tasks, such as inventory management, traffic signal optimization, network optimization, resource allocation, and robotics, are represented by structured observations with categorical or mixed data types.
-These tasks can benefit significantly from deployment and training on edge devices due to resource constraints.
-Moreover, interpretability is crucial in these applications for regulatory reasons and for trust in the decision-making process.
-Current neural network (NN) based solutions struggle with interpretability, handling categorical data, and supporting light implementations suitable for low-compute devices.
-Gradient Boosting Trees (GBT) is a powerful ensemble method extensively used in supervised learning due to its simplicity, accuracy, interpretability, and natural handling of structured and categorical data.
-Frameworks such as XGBoost [7], LightGBM [20], and CatBoost [36] have become integral in applications spanning finance [49], healthcare [54, 27, 43], and competitive data science [6].
-Despite their successes, GBT has seen limited application in RL.
-This is primarily because traditional GBT libraries are designed for static datasets with predefined labels, contrasting with the dynamic nature of RL.
-The distribution shift in both input (state) and output (reward) poses significant challenges for the direct application of GBT in RL.
-Moreover, there is a notable lack of benchmarks or environments tailored for structured data, further hindering progress in this area.
-In this paper, we introduce Gradient Boosting Reinforcement Learning (GBRL), a GBT framework tailored for RL. Our contributions are:
-  1. 1.
-    GBT for RL.
-We demonstrate the viability and potential of GBT as function approximators in RL.
-We present GBT-based implementations of PPO, A2C, and AWR, and show that GBRL is competitive with NNs across a range of environments.
-In addition, similarly to supervised learning, GBRL outperforms NNs on categorical tasks (see Figure 1).
-  2. 2.
-    Tree-based Actor-Critic architecture.
-Inspired by shared architectures in NN-based actor-critic (AC), we introduce a GBT-based AC architecture.
-This reduces the memory and computational requirements by sharing a common ensemble structure for both the policy and value.
-This approach significantly reduces runtime compared to existing GBT frameworks, thus removing the barrier to solving complex, high-dimensional RL tasks with millions of interactions.
-  3. 3.
-    Modern GBT-based RL library  [^1] The GBRL core library is available at https://github.com/NVlabs/gbrl.
-The actor-critic implementations based on GBRL are available at https://github.com/NVlabs/gbrl_sb3..
-We provide a CUDA-based [33] hardware-accelerated GBT framework optimized for RL.
-GBRL is designed to work as part of a broader system and seamlessly integrates with popular repositories such as Stable-baselines3 [39].
-This new tool offers practitioners a powerful option for exploring GBT in RL settings.
-Figure 1: PPO GBRL vs PPO NN.
-Aggregated mean and standard deviation of the normalized average reward for the final 100 episodes.
-Rewards were normalized as: $\text{reward}_{\text{norm}}=\frac{\text{reward}}{\text{reward}_{\text{max}\{\text{NN},\text{GBRL}\}}}$ per environment and then aggregated across each domain.
+Many real-world decision-making tasks involve structured observations, where data can be organized in a tabular format and follow predefined organizational patterns.
+Unlike unstructured data (such as images or raw sensor data), structured observations often include heterogeneous features, both numerical and categorical.
+These features carry direct semantic meaning that can be used for prediction without complex feature extraction.
+Such structured observations closely resemble tabular datasets commonly encountered in supervised learning.
+Notable examples where structured observations are crucial include recommendation systems, healthcare diagnostics, digital advertising, fraud detection, weather prediction, and financial trading.
+In these domains, a key challenge lies in handling structured observations, where the importance of different components varies dynamically with the agent’s state and task.
+Neural network (NN)-based solutions struggle with such data, requiring sophisticated architectures and preprocessing techniques to capture structural relationships (Zabërgja et al., 2024; Kadra et al., 2021; Arik & Pfister, 2021; Gorishniy et al., 2021; Hollmann et al., 2023).
+This often results in complex models that sacrifice sample efficiency and generalization capability.
+Gradient boosting trees (GBT) offer a promising alternative through their natural handling of structured data.
+Their success in supervised learning stems from an iterative ensemble-building process, in which each decision tree refines predictions by partitioning the input space along individual features.
+This approach has made GBT frameworks such as XGBoost (Chen & Guestrin, 2016), LightGBM (Ke et al., 2017), and CatBoost (Prokhorenkova et al., 2018) integral in domains such as finance (Tian et al., 2020), healthcare (Wassan et al., 2022; Ma et al., 2022; Seto et al., 2022), and competitive data science (Chen, 2023).
+This approach creates an ensemble of piecewise constant functions that excel at capturing non-smooth patterns and abrupt transitions in data (Jeffares et al., 2024; Beyazit et al., 2023; Grinsztajn et al., 2022).
+These properties are particularly valuable in RL, where state-action relationships often exhibit sudden changes, for instance, when transitioning between different phases of a task or when certain state variables cross critical thresholds.
+Such discontinuous relationships are common in structured RL tasks, from inventory management where actions change discretely with stock levels, to game scenarios where strategic decisions shift abruptly based on state conditions.
+Despite these potential benefits, adapting GBT to RL presents significant challenges.
+Traditional GBT frameworks are designed for static datasets with stable feature distributions and predefined loss functions, where the focus is on optimizing batch training efficiency.
+This contrasts sharply with RL’s dynamic nature, where state distributions evolve during training, data is collected iteratively, and rewards are often delayed.
+While traditional GBT libraries can technically be adapted to RL using custom loss functions and incremental learning, our experiments with CatBoost and XGBoost (Section 4.3) show that this approach fails to scale due to their supervised learning-oriented design.
+We address these challenges through Gradient Boosting Reinforcement Learning (GBRL), a framework that bridges the gap between GBT’s strengths and RL’s unique requirements. Our contributions are:
+1.
+New framework for GBT in RL: We establish GBT as a viable function approximator for RL by devising popular algorithms—PPO, A2C, and AWR—with GBT backends.
+In popular environments, our framework demonstrates competitive performance against NNs while showing particular strengths on structured tasks.
+2.
+Improved robustness: We demonstrate that GBT’s inherent strengths translate effectively to RL tasks.
+Through extensive dedicated experiments, we show superior robustness to out-of-distribution scenarios, resilience to spurious correlations, and better handling of state-space perturbations – critical advantages over NNs for real-world applications.
+3.
+Practical implementation: We provide a CUDA-accelerated (NVIDIA, 2025) implementation that seamlessly integrates with existing RL libraries like Stable-baselines3 (Raffin et al., 2021) [^1] The GBRL core library is available at https://github.com/NVlabs/gbrl.
+[^2] Actor-Critic implementations integrated within Stable-baselines3, are available at https://github.com/NVlabs/gbrl_sb3.
+Our shared Actor-Critic design achieves state-of-the-art performance while significantly reducing both memory and computational requirements on modern hardware.
 2 Related Work
-Gradient boosted trees.
-Recent advances have extended GBT’s capabilities beyond traditional regression and classification.
-In ranking problems, GBT has been used to directly optimize ranking metrics, as demonstrated by frameworks like StochasticRank [51] and recent advancements explored in Lyzhin et al.
-[26].
-Additionally, GBT offer probabilistic predictions through frameworks like NGBoost [11], enabling uncertainty quantification [28].
-The connection between GBT and Gaussian Processes [52, 45] offers further possibilities for uncertainty-aware modeling.
-Recently, Ivanov and Prokhorenkova [18] modeled graph-structured data by combining GBT with graph neural networks.
+GBT Beyond Supervised Learning.
+Recent advances have extended the capabilities of GBT beyond traditional regression and classification.
+In ranking problems, GBT has been used to directly optimize ranking metrics (Lyzhin et al., 2023), as demonstrated by frameworks such as StochasticRank (Ustimenko & Prokhorenkova, 2020).
+Furthermore, GBT offers probabilistic predictions through frameworks such as NGBoost (Duan et al., 2020), allowing uncertainty quantification (Malinin et al., 2021).
+The connection between GBT and Gaussian processes (Ustimenko et al., 2023; Sigrist, 2022) offers further possibilities for uncertainty-aware modeling.
+Recently, Ivanov & Prokhorenkova (2021) modeled graph-structured data by combining GBT with graph NNs.
 Despite their versatility, applying GBT in RL remains a relatively less explored area.
-Several works have employed GBT as a function approximator within off-policy RL methods, including its use in Q-learning [1] and in bandit settings to learn inverse propensity scores [24].
+Several works have employed GBT as a function approximator within off-policy RL methods, including its use in Q-learning (Abel et al., 2016) and in bandit settings to learn inverse propensity scores (London et al., 2023).
 Recently, Brukhim et al.
-[5] proposed a boosting framework for RL where a base policy class is incrementally enhanced using linear combinations and nonlinear transformations.
-However, these previous works have not yet demonstrated the scalability and effectiveness in complex, high-dimensional RL environments requiring extensive interactions.
-In this work, we show how to adapt the framework of GBT to successfully solve large-scale RL problems.
-Interpretability.
-Due to the inherent non-linearities, NNs are challenging to interpret and require sophisticated methods to do so.
-Interpreting NNs often involves either approximation with simpler models such as decision trees or using gradient-based techniques, which require additional forward and backward passes [16, 9, 44, 3, 37].
-On the other hand, interpretability methods for GBT can take advantage of the structure of a decision tree for high speed, efficiency, and accuracy [25, 10].
-Structured and categorical data.
-Previous work in RL has predominantly focused on using NNs due to their ability to capture complex patterns in high-dimensional data.
-Techniques such as Q-learning and AC methods have advanced significantly, demonstrating success in tasks involving raw sensory inputs like images, text, and audio.
-However, NNs that perform well on structured and categorical data typically have very specialized architectures [19, 46, 14, 2] and are not standard multi-layer perceptrons (MLPs) that are often used in many RL tasks and algorithms [34].
-Even with these specialized architectures, Gradient Boosting Trees (GBT) often perform equally or better on structured and categorical datasets [19, 31, 14, 15].
+(2022) proposed a boosting framework for RL where a base policy class is incrementally enhanced using linear combinations and nonlinear transformations, forming a 2-layer NN.
+However, these previous works have not demonstrated scalability and effectiveness in complex, high-dimensional RL environments that require extensive interactions.
+In this work, we show how to adapt the framework of GBT to successfully solve large-scale RL problems for the first time.
+Tabular Data.
+Previous work in RL has focused predominantly on the use of NNs due to their ability to capture complex patterns in high-dimensional data.
+Techniques such as Q-learning and Actor-Critic methods have advanced significantly, demonstrating success in tasks involving raw sensory inputs like images, text, and audio.
+However, NNs that perform well on tabular data typically have very specialized architectures (Katzir et al., 2021; Somepalli et al., 2021; Gorishniy et al., 2021; Arik & Pfister, 2021) and are different from the standard multi-layer perceptrons that are often used in RL tasks and algorithms (Ota et al., 2024).
+Even with specialized architectures, GBT often performs equally or better on tabular datasets (Borisov et al., 2021; McElfresh et al., 2023; Grinsztajn et al., 2022; Shwartz-Ziv & Armon, 2022).
+GBT’s success can be attributed to a key difference in inductive biases compared to NNs.
+Jeffares et al.
+(2024) have shown that the kernel representations induced by GBTs are bounded and often behave more predictably on irregular or out-of-distribution inputs.
+In contrast, NN-tangent kernels can be unbounded and vary significantly on test points far from the training distribution.
 Policy Optimization through Functional Gradient Ascent.
-In this approach, the policy is parameterized by a growing linear combination of functions [29].
-Each linear addition represents the functional gradient with respect to current parameters.
-Kersting and Driessens [21] demonstrated the direct optimization of policies using the policy gradient theorem [48].
-Similarly, Scherrer and Geist [41] proposed a functional gradient ascent approach as a local policy search algorithm.
-While these works lay theoretical groundwork, practical results on complex, high-dimensional RL environments have not been shown.
-To adapt GBT’s to RL, we leverage the framework of functional gradient ascent.
-This combination enables a seamless integration of GBRL directly into existing RL optimization packages, such as Stable-baselines [39].
-3 Preliminaries
-We begin by introducing Markov Decision Processes (MDPs) and the AC schema. Then, we introduce GBT. In the following section, we show how to combine both of these paradigms into GBRL.
-3.1 Markov Decision Process
-We consider a fully observable infinite-horizon Markov decision process (MDP) characterized by the tuple $(\mathcal{S},\mathcal{A},P,\mathcal{R})$.
-At each step, the agent observes a state $\operatorname{\mathbf{s}}\in\mathcal{S}$ and samples an action $\operatorname{\mathbf{a}}\in\mathcal{A}$ from its policy $\pi(\operatorname{\mathbf{s}},\operatorname{\mathbf{a}})$.
-Performing the action causes the system to transition to a new state $\operatorname{\mathbf{s}}^{\prime}$ based on the transition probabilities $P(\operatorname{\mathbf{s}}^{\prime}|\operatorname{\mathbf{s}},\operatorname{\mathbf{a}})$, and the agent receives a reward $\operatorname{\mathbf{r}}\sim\mathcal{R}(\operatorname{\mathbf{s}},\operatorname{\mathbf{a}})$.
-The objective is to find an optimal policy $\pi^{*}$ that maximizes the expected discounted reward $J(\pi)=\mathbb{E}[\sum_{t=0}^{\infty}\gamma^{t}\operatorname{\mathbf{r}}_{t}]$, with a discount factor $\gamma\in[0,1)$.
-The action-value function $Q_{\pi}(\operatorname{\mathbf{s}},\operatorname{\mathbf{a}}):=\mathbb{E}_{\pi}[\sum_{t^{\prime}=0}^{\infty}\gamma^{t^{\prime}}\mathcal{R}(\operatorname{\mathbf{s}}_{t+t^{\prime}},\operatorname{\mathbf{a}}_{t+t^{\prime}})|\operatorname{\mathbf{s}}_{t}=\operatorname{\mathbf{s}},\operatorname{\mathbf{a}}_{t}=\operatorname{\mathbf{a}}]$ estimates the expected returns of performing action $\operatorname{\mathbf{a}}$ in state $\operatorname{\mathbf{s}}$ and then acting according to $\pi$.
-Additionally, the value function $V_{\pi}(\operatorname{\mathbf{s}}):=\mathbb{E}_{\pi}[\sum_{t^{\prime}=0}^{\infty}\gamma^{t^{\prime}}\mathcal{R}(\operatorname{\mathbf{s}}_{t+t^{\prime}},\operatorname{\mathbf{a}}_{t+t^{\prime}})|\operatorname{\mathbf{s}}_{t}=\operatorname{\mathbf{s}}]$, predicts the expected return starting from state $\operatorname{\mathbf{s}}$ and acting according to $\pi$.
-Finally, the advantage function $A_{\pi}(\operatorname{\mathbf{s}},\operatorname{\mathbf{a}}):=Q_{\pi}(\operatorname{\mathbf{s}},\operatorname{\mathbf{a}})-V_{\pi}(\operatorname{\mathbf{s}})$, indicates the expected relative benefit of performing action $\operatorname{\mathbf{a}}$ over acting according to $\pi$.
-3.2 Actor-Critic Reinforcement Learning
-Actor-critic methods are a common method to solve the objective $J(\pi)$.
-They learn both the policy and value.
-In the GBRL framework, we extend three common AC algorithms to support GBT-based function approximators.
-A2C [32] is a synchronous, on-policy AC algorithm designed to improve learning stability.
-The critic learns a value function, $V(\operatorname{\mathbf{s}})$, used to estimate the advantage.
-This advantage is incorporated into the policy gradient updates, reducing variance and leading to smoother learning.
-The policy is updated using the following gradient: $\nabla_{\theta}J(\pi_{\theta})=\mathbb{E}[\nabla_{\theta}\log\pi_{\theta}(\operatorname{\mathbf{a}}|\operatorname{\mathbf{s}})A(\operatorname{\mathbf{s}},\operatorname{\mathbf{a}})]$.
-PPO [42] extends A2C by improving stability.
-This is achieved through constrained policy update steps using a clipped surrogate objective.
-This prevents drastic policy changes and leads to smoother learning.
-To achieve this, PPO solves the following objective: $\nabla_{\theta}J(\pi_{\theta})=\mathbb{E}[\nabla_{\theta}\text{clip}(\frac{\log\pi_{\theta}(\operatorname{\mathbf{a}}|\operatorname{\mathbf{s}})}{\log\pi_{\theta_{\text{old}}}(\operatorname{\mathbf{a}}|\operatorname{\mathbf{s}})},1-\epsilon,1+\epsilon)A(\operatorname{\mathbf{s}},\operatorname{\mathbf{a}})]$.
-Additionally, PPO enhances sample efficiency by performing multiple optimization steps on each collected rollout.
-AWR [35] is an off-policy AC algorithm.
-Provided a dataset $\mathcal{D}$, AWR updates both the policy and the value through supervised learning.
-This dataset can be pre-defined and fixed (offline), or continually updated using the agents experience (replay buffer).
-At each training iteration $k$, AWR solves the following two regression problems:
-$V_{k}$ $=\operatorname*{arg\,min}_{V}\mathbb{E}_{\operatorname{\mathbf{s}},\operatorname{\mathbf{a}}\sim\mathcal{D}}[\|G(\operatorname{\mathbf{s}},\operatorname{\mathbf{a}})-V(\operatorname{\mathbf{s}})\|_{2}^{2}]\,,\,\pi_{k+1}=\operatorname*{arg\,max}_{\pi}\mathbb{E}_{\operatorname{\mathbf{s}},\operatorname{\mathbf{a}}\sim\mathcal{D}}[\log\pi(\operatorname{\mathbf{a}}|\operatorname{\mathbf{s}})\exp(\frac{1}{\beta}A_{k}(\operatorname{\mathbf{s}},\operatorname{\mathbf{a}}))]\,,$
-where $G(\operatorname{\mathbf{s}},\operatorname{\mathbf{a}})$ represents the monte-carlo estimate or TD($\lambda$) of the expected return [47].
-3.3 Gradient Boosting Trees as Functional Gradient Descent
-Gradient boosting trees (GBT) [12] are a non-parametric machine learning technique that combines decision tree ensembles with functional gradient descent [30].
-GBT iteratively minimizes the expected loss $L(F(\mathbf{x}))=\mathbb{E}_{\mathbf{x},\mathbf{y}}[L(\mathbf{y},F(\mathbf{x}))]$ over a dataset $D=\{(\mathbf{x}_{i},\mathbf{y}_{i})\}_{i=1}^{N}$.
-A GBT model, $F_{K}$, predicts outputs using $K$ additive trees as follows:
-$F_{K}(\mathbf{x}_{i})=F_{0}+\sum_{k=1}^{K}\epsilon h_{k}(\mathbf{x}_{i})\,,$ (1)
-where $\epsilon$ is the learning rate, $F_{0}$ is the base learner, and each $h_{k}$ is an independent regression tree partitioning the feature space.
-In the context of functional gradient descent, the objective is to minimize the expected loss $L(F(\mathbf{x}))=\mathbb{E}_{\mathbf{x},\mathbf{y}}[L(\mathbf{y},F(\mathbf{x}))]$ with respect to the functional $F$.
-Here, a functional $F:\mathcal{H}\rightarrow\mathbb{R}$ maps a function space to real numbers.
-A GBT model can be viewed as a functional $F$ that maps a linear combination of binary decision trees to outputs: $F:\text{lin}(\mathcal{H})\rightarrow\mathbb{R}^{D}$, where $\mathcal{H}$ is the decision tree function class.
-We start with an initial model, $F_{0}$, and iteratively add trees to $F$ to minimize the expected loss.
-Similar to parametric gradient descent, at each iteration $k$, we minimize the loss by taking a step in the direction of the functional gradient $g^{i}_{k}:=\nabla_{F_{k-1}}L(\mathbf{y}_{i},F_{k-1}(\mathbf{x}_{i}))$.
-However, we are constrained to gradient directions within $\mathcal{H}$.
-Thus, we project the gradient $g_{k}$ into a decision tree by solving:
-$h_{k}=\arg\min_{h}\|-\epsilon g_{k}-h(\mathbf{x})\|_{2}^{2}\,.$ (2)
+Kersting & Driessens (2008) introduced Non-Parametric Policy Gradients (NPPG).
+By combining the policy gradient theorem (Sutton et al., 1999) with functional gradient ascent, NPPG directly optimizes the policy.
+Although NPPG sets the foundation for combining GBT with RL, it is limited to discrete action spaces and policies represented by the Gibbs distribution.
+GBRL extends the ideas laid in NPPG to general Actor-Critic algorithms and generalizes to any policy representation and objective.
+Figure 1: The GBRL framework.
+The actor’s policy and critic’s value function are parameterized by the tree ensemble $F_{k}$.
+For example, $F_{k}(\operatorname{s})=[\mu(\operatorname{s}),\sigma(\operatorname{s}),V(\operatorname{s})]$ for a Gaussian policy.
+Starting from $F_{0}$, at each training iteration, $k$, GBRL collects a rollout and computes the gradient $\nabla_{\pi_{F_{k}}}J(\pi_{F_{k}})$ with respect to the current ensemble.
+This gradient is then used to fit the next tree.
+Adding the tree to the ensemble updates it to $F_{k}(\operatorname{s})=F_{k-1}(\operatorname{s})+\epsilon h_{k}(\operatorname{s})$, where $\epsilon$ is the learning rate.
+3 RL Preliminaries
+A fully observable infinite-horizon MDP is characterized by the tuple $(\mathcal{S},\mathcal{A},P,\mathcal{R})$.
+At each step, the agent observes a state $\operatorname{s}\in\mathcal{S}$ and samples an action $\operatorname{a}\in\mathcal{A}$ from its policy $\pi(\operatorname{s},\operatorname{a})$.
+The action causes the system to transition to a new state $\operatorname{s}^{\prime}$ based on the transition probabilities $P(\operatorname{s}^{\prime}|\operatorname{s},\operatorname{a})$, and produce a reward $\operatorname{r}\sim\mathcal{R}(\operatorname{s},\operatorname{a})$.
+The agent’s objective is to maximize the expected discounted reward $J(\pi)=\mathbb{E}[\sum_{t=0}^{\infty}\gamma^{t}\operatorname{r}_{t}]$, with a discount factor $\gamma\in[0,1)$.
+The state-action value function $Q_{\pi}(\operatorname{s},\operatorname{a}):=\mathbb{E}_{\pi}[\sum_{t=0}^{\infty}\gamma^{t}\operatorname{r}_{t}|\operatorname{s}_{0}=\operatorname{s},\operatorname{a}_{0}=\operatorname{a}]$ estimates the expected returns of performing action $\operatorname{a}$ in state $\operatorname{s}$ and then acting according to $\pi$.
+Additionally, the value function $V_{\pi}(\operatorname{s}):=\mathbb{E}_{\pi}[\sum_{t=0}^{\infty}\gamma^{t}\operatorname{r}_{t}|\operatorname{s}_{0}=\operatorname{s}]$, predicts the expected return starting from state $\operatorname{s}$ and acting according to $\pi$.
+Finally, the advantage function $A_{\pi}(\operatorname{s},\operatorname{a}):=Q_{\pi}(\operatorname{s},\operatorname{a})-V_{\pi}(\operatorname{s})$, indicates the expected relative benefit of performing action $\operatorname{a}$ over acting according to $\pi$.
+Actor-Critic methods are a common approach in RL.
+They simultaneously learn a policy and a value function to improve the efficiency and stability of training.
+For conciseness, in the body of the paper we focus on Proximal Policy Optimization (PPO) (Schulman et al., 2017) which jointly trains an actor and critic.
+In the supplementary material, we also showcase the implementation and results for A2C (Mnih et al., 2016) and AWR (Peng et al., 2019).
 4 Gradient Boosting Reinforcement Learning
-Figure 2: The GBRL framework.
-The actor’s policy and critic’s value function are parameterized by $\theta_{k}$.
-For example, $\theta_{k}=[\mu(\operatorname{\mathbf{s}}),\sigma(\operatorname{\mathbf{s}}),V(\operatorname{\mathbf{s}})]$ for a Gaussian policy.
-$\theta_{k}$ is calculated by summing all the outputs of trees in the ensemble.
-Starting from $\theta_{0}$, at each training iteration, GBRL collects a rollout and computes the gradient $\nabla_{\theta_{0}}J$.
-This gradient is then used to fit the next tree added to the ensemble, which is updated to $\theta_{1}$.
-This process repeats with each iteration fitting a new tree, refining the parameterization, and expanding the ensemble towards $\theta_{k}\approx\theta_{0}+\epsilon\sum^{k-1}_{m=0}\nabla_{\theta_{m}}J$, an approximated scaled sum of gradients with respect to past parametrizations.
-In this work, we extend the framework of GBT to support AC algorithms in the task of RL.
-The objective in RL is to optimize the return $J$, the cumulative reward an agent receives.
-Unlike in supervised learning, the target predictions are unknown a priori.
-RL agents learn through trial and error.
-Good actions are reinforced by taking a step in the direction of the gradient $\nabla_{\pi}J$.
-This formulation aligns perfectly with functional gradient ascent; thus, in GBRL, we optimize the objective directly over the decision tree function class.
-This is achieved by iteratively growing the ensemble of trees $\{h_{i}\}$.
-The ensemble outputs $\theta$, representing AC parameters such as the policy $\pi$ and the value function.
-For example, $\theta=[\mu(\operatorname{\mathbf{s}}),\sigma(\operatorname{\mathbf{s}}),V(\operatorname{\mathbf{s}})]$ for a Gaussian policy.
-At each iteration, a new tree $h_{k}$, constructed to minimize the distance to $\nabla_{\theta_{k-1}}J$, is added to the ensemble.
-Here, The resulting method is an application of GBT as a functional gradient optimizer $\theta_{k}\approx\theta_{0}+\epsilon\sum^{k-1}_{m=0}\nabla_{\theta_{m}}J$.
-However, RL presents unique challenges for GBT.
-RL involves a nonstationary state distribution and inherent online learning, causing gradients to vary in magnitude and direction.
-Large gradients in unfavorable directions risk destabilizing training or leading to catastrophic forgetting.
-Moreover, feedback in RL is provided through interactions with the environment and is not available a priori.
-This contrasts with supervised learning settings, where gradients decrease with boosting iterations, and targets are predefined.
-As a result, many of the key features that traditional GBT libraries rely on are not suitable.
-For example, GOSS [20], categorical feature encoding [36], early-stopping signals, pruning methods [53], and strategies to tackle online learning [55].
-To address these challenges, we employ appropriate tools from the NN and GBT literature, such as batch learning [13, 40] to update the ensemble.
-At each boosting iteration, we fit a decision tree on a random batch sampled with replacement from the experience buffer.
-This approach helps handle non-stationary distributions and improve stability by focusing on different parts of the state space, allowing beneficial gradient directions to accumulate and minimizing the impact of detrimental ones.
-Additionally, GBRL fits gradients directly to optimize objectives, whereas traditional GBT methods require targets and need workarounds to utilize gradients effectively.
-A common theme in AC algorithms is to utilize a shared approximation for the actor and the critic.
-We adopt this approach in GBRL, constructing trees where each leaf provides two predictions.
+Deep RL with NNs faces significant challenges when dealing with structured observations containing both numerical and categorical features.
+Similarly to tabular datasets in supervised learning, these observations consist of heterogeneous features representing fundamentally different types of information, such as numerical measurements alongside categorical attributes describing system states or task conditions.
+These features create highly irregular patterns, as they may independently affect the environment based on their context, scale, or statistical properties.
+NNs struggle to model such patterns effectively (Beyazit et al., 2023; Jeffares et al., 2024), and their performance heavily depends on preprocessing methods (Hancock & Khoshgoftaar, 2020; Gorishniy et al., 2022).
+In contrast, GBT naturally handles irregular patterns and structured feature spaces efficiently.
+To leverage this strength, we present GBRL, an adaptation of GBT to RL tasks with such representations that frequently arise in real-world problems.
+4.1 GBT as Functional Gradient Descent
+GBT (Friedman, 2001) combines decision tree ensembles with functional gradient descent to learn complex nonlinear functions (Mason et al., 1999).
+Given a loss function $L$ and a dataset, $D=\{(\mathbf{x}_{i},\mathbf{y}_{i})\}_{i=1}^{N}$, where $\mathbf{x}_{i}\in\mathbb{R}^{M}$ and $\mathbf{y}_{i}\in\mathbb{R}^{D}$, GBT aims to learn a function $F$, mapping inputs $\mathbf{x}$ to outputs $\mathbf{y}$.
+Through functional gradient descent, GBT minimizes the expected loss $\mathbb{E}_{\mathbf{x},\mathbf{y}}[L(\mathbf{y},F(\mathbf{x}))]$ with respect to the learned function $F$.
+Unlike parametric gradient descent, which updates parameters, functional gradient descent updates a learned function directly.
+Starting with an initial function $F_{0}$.
+At each iteration $k$, GBT refines $F_{k}$, by taking a step in the negative direction of the gradient $g_{k}:=\nabla_{F_{k}}L(\mathbf{y},F_{k}(\mathbf{x}))$.
+The negative gradient is projected onto the space of possible trees $h\in\mathcal{H}$ by constructing a new binary decision tree that minimizes the following objective:
+$h_{k}=\arg\min_{h\in\mathcal{H}}\|-g_{k}-h(\mathbf{x})\|_{2}^{2}\,.$ (1)
+Through this process, GBT minimizes the expected loss, resulting in the additive model:
+$F_{K}(\mathbf{x})=F_{0}+\sum_{k=1}^{K-1}\epsilon h_{k}(\mathbf{x})\,,$ (2)
+where, $\epsilon$ is the learning rate controlling the step size in each iteration.
+While standard GBT frameworks excel at handling structured data in supervised tasks, they require a tailored solution for RL’s iterative, dynamic, and multi-objective nature.
+4.2 The GBRL Framework
+Library Incremental Learning (sec) Standard Learning (sec)
+CatBoost $36.17\pm 2.77$ $1.15\pm 0.03$
+XGBoost $67.11\pm 0.70$ $2.26\pm 3.70$
+GBRL $\mathbf{5.47}\pm\mathbf{1.19}$ N/A
+Table 1: Incremental Learning Speed in GBT Libraries.
+Training times (seconds) for 1000 boosting iterations on random batches of 128 samples with 20 input features and 3 targets.
+While CatBoost and XGBoost train efficiently in standard mode using the same batch for all iterations, their training time increases significantly in incremental learning, where each batch receives only one boosting iteration.
+GBRL, designed explicitly for incremental updates, remains much faster.
+In the GBRL framework (Figure 1) we adapt GBTs to handle the challenges of RL.
+To do so we reinterpret $F_{K}(\operatorname{s})$ as the parameterization of the policy and value function, where the observed state $\operatorname{s}\in\mathcal{S}$ is the input to the ensemble.
+Specifically, at each iteration $k$, we compute the policy gradient (Sutton et al., 1999).
+As opposed to the original formulation, where the gradient is calculated with respect to parameter space, here we calculate it in function space.
+Consequently, we express the policy gradient as:
+$\nabla_{F_{k}}J(\pi_{F_{k}})=\mathbb{E}_{\pi_{F_{k}}}[\nabla_{F_{k}}\log\pi_{F_{k}}(\operatorname{a}|\operatorname{s})A(\operatorname{s},\operatorname{a})],$ (3)
+where $F_{k}$ is the current ensemble-based parametrization.
+Then, the policy gradient is projected onto a new tree and added to the ensemble:
+$h_{k}=\arg\min_{h\in\mathcal{H}}\|\nabla_{F_{k}}J(\pi_{F_{k}})-h(\operatorname{s})\|_{2}^{2}\,.$ (4)
+The result is an application of GBT as a functional gradient optimizer, updating both actor and critic incrementally:
+$F_{K}(\operatorname{s})=F_{0}+\sum_{k=1}^{K-1}\epsilon h_{k}(\operatorname{s})\approx F_{0}+\sum^{K-1}_{k=1}\epsilon\nabla_{F_{k}}J(\pi_{F_{k}}).$ (5)
+Gradient-based Approach.
+As seen in Equation 5, GBRL is an online learning gradient-based optimizer.
+In each step, a new tree is constructed to minimize the sampled loss.
+This learning scheme is identical to the common policy gradient methods (Mnih et al., 2016).
+By leveraging optimization frameworks for gradient computation, such as PyTorch (Paszke et al., 2019), GBRL can be integrated with most Actor-Critic algorithms and implemented within existing RL libraries.
+In contrast, traditional GBT frameworks are designed for offline, static datasets and require significant customizations and workarounds to function in RL.
+Figure 2: GBT library comparison, Cartpole. CatBoost and XGBoost are intractable in RL. CatBoost’s lack of GPU support for custom losses leads to low FPS and early termination.
+Figure 3: Shared Actor-Critic, MiniGrid.
+The shared tree structure significantly increases efficiency, without impacting the score.
+Aggregated results over three tasks are shown here, full per-task curves are available in Appendix (Figure 20)
+Shared Actor-Critic Architecture.
+Many RL algorithms operate with a shared Actor-Critic structure.
+This structure helps avoid overfitting and improves overall convergence speeds (Henderson et al., 2018; Andrychowicz et al., 2020).
+We adopt this approach in GBRL such that each leaf produces two predictions.
 GBRL predicts both the policy (distribution over actions) and the value estimate.
-The internal structure of the tree is shared, providing a single feature representation for both objectives and significantly reducing memory and computational bottlenecks.
-Accordingly, in GBRL we apply differentiated learning rates to the policy and value outputs during prediction, effectively optimizing distinct objectives within this shared structure.
-We present the full algorithm in Algorithm 1 and diagram in Figure 2.
-Algorithm 1 Gradient Boosting for Reinforcement Learning (GBRL)
-1:Initialize: $\theta_{0}$, $\epsilon_{\text{actor}}$, $\epsilon_{\text{critic}}$, experience buffer $\mathcal{B}$, total training iterations $K$, number of updates $U$, batch size $N$, $k\leftarrow 1$
-2:while $k<K$ do
-3:     Collect trajectory $\tau^{(k)}=(\operatorname{\mathbf{s}}_{0},\operatorname{\mathbf{a}}_{0},\ldots,\operatorname{\mathbf{s}}_{T},\operatorname{\mathbf{a}}_{T})^{(k)}$ and rewards $(\operatorname{\mathbf{r}}_{0},\ldots,\operatorname{\mathbf{r}}_{T})^{(k)}$ using $\pi_{\theta_{k-1}}$
-4:     Add trajectory $\tau^{(k)}$ and rewards to the experience buffer $\mathcal{B}$
-5:     for each update $u=1$ to $U$ do
-6:         Sample a batch from the experience buffer $\mathcal{B}$
-7:         Compute gradients $g$ according to AC algorithm (e.g., PPO, A2C, AWR)
-8:         Construct dataset $D=\{(\operatorname{\mathbf{s}}_{n},g_{n})\}_{n=0}^{N}$ and fit a decision tree $h_{k}$
-9:         for each dimension $d=0$ to $D$ do
-10:              if $0\leq d<D$ then
-11:                  Update $\theta^{(d)}_{k}=\theta^{(d)}_{k-1}+\epsilon_{\text{actor}}h_{k}(\operatorname{\mathbf{s}})$
-12:              else
-13:                  Update $\theta^{(d)}_{k}=\theta^{(d)}_{k-1}+\epsilon_{\text{critic}}h_{k}(\operatorname{\mathbf{s}})$
-14:         $k\leftarrow k+1$
-15:Output: AC parameters $\theta^{(d)}_{K}(\operatorname{\mathbf{s}})$ for $d=0,1,\ldots,D$
-5 Experiments
-Our experiments aim to answer the following questions:
-  1. 1.
-    GBT as RL Function Approximator: Can GBT-based AC algorithms effectively solve complex high-dimensional RL tasks?
-  2. 2.
-    Comparison to NNs: How does GBRL compare with NN-based training in various RL algorithms?
-  3. 3.
-    Benefits in Categorical Domains: Do the benefits of GBT in supervised learning transfer to the realm of RL?
-  4. 4.
-    Comparison to Traditional GBT libraries: Can we use traditional GBT libraries instead of GBRL for RL tasks?
-  5. 5.
-    Evaluating the shared AC architecture: How does sharing the tree structure between the actor and the critic impact performance?
-We implemented GBT-based versions of A2C, PPO, and AWR within Stable Baselines3.
-We refer to our implementations as PPO GBRL, A2C GBRL, and AWR GBRL.
-We evaluated GBRL against the equivalent NN implementations.
-Where available, we utilize hyperparameters from RL Baselines3 Zoo [38]; otherwise, we optimize the hyperparameters for specific environments.
-The AWR NN implementation is based on the original paper [35].
-We conducted experiments on a range of RL domains.
-We test classic control tasks, high-dimensional vectorized problems, and finally categorical tasks.
-We use 5 random seeds per experiment on a single NVIDIA V100-32GB GPU.
-We present the cummulative non-discounted reward, averaged across the last 100 episodes.
-We normalize the plots for simple visual comparison between GBRL and the corresponding NN implementations.
-The normalized score is computed as $\text{score}_{\text{norm}}=\frac{\text{score}_{\text{GBRL}}-\text{score}_{\text{NN}}}{\text{score}_{\text{max}\{\text{NN},\text{GBRL}\}}-\text{score}_{\text{min}\{\text{NN},\text{GBRL}\}}}$.
-We provide the full learning curves, implementation details, compute resources, un-normalized numerical results, and hyperparameters in the supplementary material.
-Classic Enviroments.
-We evaluate GBRL’s ability to solve classic RL tasks using Continuous-Control and Box2D environments, provided via Gym [50].
-We trained agents for 1M steps (1.5M for LunarLander-v2) and provide the results in Figure 3.
-For exact values, refer to Table 2.
-Considering the algorithmic objective, we observe that GBRL and NN present similar performance when optimized using PPO.
-In contrast, the other methods demonstrate inconclusive results.
-In certain environments, such as MountainCar, GBRL outperforms NN with all AC methods.
-On the other hand, in Pendulum NN is better.
-Figure 3: Continuous-Control and Box2D environments. Normalized comparison between GBRL and NN. PPO, the best performing method, shows similar performance with GBRL and NN function classes.
-High-Dimensional Vectorized Environments.
-The decision-tree function class operates on individual features at each step.
-Consequently, this function class is not well-suited for handling pixel-based representations, which require more complex feature interactions.
-Therefore, we evaluted GBRL in the Football [22] and Atari RAM [4] domains.
-These offer high-dimensional vectorized representations.
-We trained agents in both environments for 10 million timesteps.
-The complete results are reported in Tables 3 and 4 and illustrated in Figure 4.
-The results portray the following phenomenon.
-While both tasks may seem similar, there is a distinct difference.
-The features in the football domain are manually constructed and represent identifiable information, such as the location of the ball and the players.
-However, the Atari RAM domain provides a flattened view of the system RAM, which is unstructured.
-At their core, binary decision trees are if-else clauses.
-This function class is naturally suited to work with structured data.
-These insights are emphasized in the football domain.
-Here, PPO GBRL greatly outperforms PPO NN across most environments and exhibits equivalent performance on the rest.
-In addition, as Atari RAM is unstructured, we observe that, as can be expected, in most cases GBRL underperforms NN, except for AWR.
-However, AWR NN underperformed considerably compared to the other NN implementations.
-(a) Football Academy environments.
-(b) Atari-ramNoFrameskip-v4 environments.
-Figure 4: High-Dimensional Vectorized Environments. GBRL outperforms NN on the structured Football domain using PPO. NN outperforms on unstructured tasks, such as Atari RAM.
-Categorical Environments.
-The football experiment suggests that GBRL outperforms when assigned structured data.
-Here, we focus on categorical environments.
-This is a regime where GBT excels in supervised learning.
-In these experiments, we evaluated the MiniGrid domain [8].
-It consists of 2D grid worlds with goal-oriented tasks that require object interaction.
-We trained in PutNear, FourRooms, and Fetch tasks for 10M timesteps, matching the reported PPO NN in RL Baselines3 Zoo.
-We trained the remaining environments for 1M timesteps.
-We give the results in Figure 5.
-For exact numbers, see Table 5.
-Figure 5: MiniGrid environments. GBRL combined with the PPO backend outperforms NN across a range of categorical environments.
-In MiniGrid, GBRL outperforms or is on-par with NN in most tasks.
-Specifically, PPO GBRL is significantly better than PPO NN.
-We observe the same trend when comparing between environments.
-These results emphasize that GBRL is a strong candidate for problems characterized by structured data, specifically when using PPO as the algorithmic backend.
-GBRL vs Traditional GBT Libraries.
-Here, we compare GBRL with Catboost and XGBoost.
-We focus on the PPO variant.
-When comparing to the standard libraries, we utilize their built-in options for incremental learning, vectorized leaves, and custom loss functions.
-As both CatBoost and XGBoost do not support differential learning rates, we used separate ensembles for the actor and the critic.
-For the comparison, we use the CartPole-v1 environment, training for 1M steps.
-The results are shown in Figure 6.
-As seen, standard GBT libraries are unable to solve RL tasks in a realistic timeframe.
-GBRL, however, efficiently solves the task while also remaining competitive with NN across a range of environments.
-Figure 6: Comparing to standard GBT libraries. CatBoost and XGBoost are intractable in RL. Specifically, CatBoost lacks GPU support for custom losses, leading to low FPS and early termination.
-Evaluating the shared AC architecture.
-Finally, we evaluated the benefits of using a shared AC architecture by training PPO GBRL on three MiniGrid environments.
-We train agents with shared and non-shared architectures for 10M timesteps and compare the score, GPU memory usage, and FPS.
-We provide the aggregated results in Figure 7, and environment-specific breakdowns in the supplementary.
-The benefit of the shared structure is clear both in terms of GPU memory consumption and computation speed.
-By sharing the tree structure, GBRL requires less than half the memory and almost triples the training FPS.
-This is achieved without any negative performance on the resulting policy, as seen in the reward plot.
-Figure 7: Shared Actor-Critic. Sharing the tree structure significantly increases training efficiency and memory, without impacting on the score.
-Result summary.
-The performance of GBRL varied across RL algorithms, but environments like MiniGrid highlight the potential advantages of using GBT in RL.
-The results suggest that GBT’s strengths in handling structured and categorical data from supervised learning can effectively transfer to the RL domain.
-Conversely, GBRL underperformed in Atari-RAM environments, indicating that certain environments, characterized by unstructured observations, are less suited for GBTs.
-The results can be explained by the findings of Grinsztajn et al.
-[15], which suggest that NNs have an inductive bias toward overly smooth solutions and that MLP-like architectures are not robust to uninformative features.
-The optimal solutions for Atari-RAM might be smoother, which could explain the better performance of NNs.
-On the other hand, McElfresh et al.
-[31] argue that GBT outperforms NNs on ‘irregular’ datasets.
-Tree-based models excel in handling irregular patterns and categorical data, aligning with GBRL’s success in environments like MiniGrid.
-Comparing different algorithmic backbones, we find PPO to be the strongest.
-PPO GBRL excelled in the MiniGrid and Football domains, and performed comparably with NN in classic control tasks.
-PPO GBRL’s success can be attributed to its alignment with GBRL’s incremental learning strategy.
-On the other hand, A2C’s single gradient update per rollout may limit its effectiveness and contribute to its underwhelming performance in many environments.
-Similarly, AWR’s design for multiple sample updates results in very large ensembles, creating a trade-off between large, slow, and memory-intensive ensembles, and lighter, less performant versions.
+The internal structure of the tree is shared, providing a single feature representation for both objectives and significantly reduces memory and computational bottlenecks.
+Similar to the standard practice, we support and use separate learning rates for the policy and value.
+This approach enables solving two distinct objectives within a single shared structure.
+4.3 Comparison to Traditional GBT Libraries
+Before diving into the comparisons with NNs, Section 5, we first evaluated our design choices against traditional GBT libraries.
+While these libraries can be adapted for RL using custom loss functions and incremental learning, experiments with CatBoost and XGBoost reveal scalability issues inherent to their supervised learning design.
+Table 1 shows these frameworks are significantly slower than GBRL, when used for incremental learning.
+Incremental learning is a fundamental property of RL, as the policy is continually changing.
+Moreover, when training on CartPole (Figure 2), larger ensemble sizes lead to significant drops in training throughput.
+In contrast, our shared Actor-Critic architecture halves memory usage and doubles throughput without compromising policy quality (Figure 3).
+Overall, GBRL offers a purpose-built solution that maintains GBT’s advantages while efficiently handling RL-specific needs.
+5 Experiments and Results
+Our experiments aim to answer two core questions:
+1. GBT as an RL function approximator: Can GBT-based AC algorithms effectively solve complex high-dimensional RL tasks? And how do they compare with NNs?
+2.
+Advantages of GBT: Building on the GBT’s success in irregular and tabular data, does its inductive bias offer similar robustness benefits in RL – specifically for out-of-distribution states, noisy inputs, and spurious correlations?
+For our experiments, we implemented a GBT-based version of PPO within Stable Baselines3 (Raffin et al., 2021).
+Where available, we use standard hyperparameters, environment-specific, and normalization wrappers according to RL Baselines3 Zoo (Raffin, 2020); otherwise, we optimize the hyperparameters for specific environments.
+For each experiment, we report the aggregated performance across five random seeds.
+Our training setup consists of a single NVIDIA V100 GPU.
+We refer the reader to the supplementary material for additional technical details, such as hyperparameters, implementation, and environment details (Appendix C), full results (Appendix D) and full learning curves (Appendix E) for all evaluated algorithms.
+5.1 Standard Environments
+Experimental Setup.
+We evaluated GBRL against NNs across three categories of RL benchmarks.
+First, we tested classic RL tasks using Classic-Control and Box2D environments from Gymnasium (Towers et al., 2024).
+These provided a baseline for comparison on standard benchmarks.
+We then evaluated both methods on more complex environments with high-dimensional vectorized representations: the Football domain (Kurach et al., 2020), where we employed the built-in ’Checkpoints’ shaped reward, and the Atari RAM domain (Bellemare et al., 2013).
+Finally, we assessed performance on categorical environments, specifically targeting the MiniGrid domain (Chevalier-Boisvert et al., 2023) – a setting where GBTs have traditionally excelled in supervised learning.
+MiniGrid provides an ideal test environment through its 2D grid-world structure featuring goal-oriented tasks with discrete object interactions.
+Results.
+We report the cumulative non-discounted reward, averaged across the last 100 episodes in Figure 4.
+For a simple visual comparison between GBRL and NNs, we report the normalized score: $\operatorname{r}_{\text{norm}}=\frac{\operatorname{r}_{\text{GBRL}}-\operatorname{r}_{\text{NN}}}{\operatorname{r}_{\text{max}\{\text{NN},\text{GBRL}\}}-\operatorname{r}_{\text{min}\{\text{NN},\text{GBRL}\}}}$.
+Figure 4: GBRL vs NN in standard environments (PPO). Aggregated mean and standard deviation of the normalized average reward for the final 100 episodes.
+Decision trees, at their core, are if-else structures that excel in handling numerical and categorical data.
+In MiniGrid, which provides such environments, GBRL significantly outperformed NN.
+For example, in tasks like PutNear, FourRooms, and Fetch, GBRL consistently achieved higher rewards, highlighting its ability to exploit the structured nature of the environment effectively.
+This performance demonstrates that GBRL is particularly well-suited for problems where data can be neatly partitioned, aligning with GBT’s inherent design strengths.
+This suggests that structured environments represent an ideal use case for GBRL.
+An additional domain characterized by structured representations is the Football domain, where we used a shaped reward.
+Here, features such as player positions and ball locations were designed to represent identifiable, interpretable information.
+In this domain, GBRL outperformed NN across most scenarios and exhibited equivalent performance in the rest.
+The structured nature of the Football domain aligns well with the strengths of GBT, effectively partitioning the feature space.
+Interestingly, while both Football and MiniGrid are structured, the Football domain is not sparse.
+Hence, GBRL’s strong performance in Football underscores its robustness and capacity to generalize well in structured, high-dimensional environments.
+Unlike Football, the Atari RAM domain offers unstructured, high-dimensional representations by exposing raw system memory states.
+These flattened views lack explicit semantic structure, making them challenging for GBRL, which relies on single-feature splits.
+As a result, GBRL underperformed compared to NN in most Atari RAM tasks.
+This result highlights an important limitation of decision-tree-based approaches.
+They struggle to model implicit feature interactions, an inherent characteristic of unstructured data.
+Finally, classic control tasks are simpler environments that rely on low-dimensional state spaces.
+In these tasks, GBRL demonstrated performance comparable to that of the NN counterpart.
+This suggests that decision tree-based models can match NNs in solving tasks where feature interactions are limited and the optimization landscape is smooth.
+Overall, these comparisons indicate that GBT can serve as a strong function approximator in RL, matching or surpassing NN’s performance in certain environments.
+5.2 Challenge Environments
+Having established GBRL’s general capabilities, we now examine scenarios designed to test inherent properties of decision trees.
+These experiments focus on challenging conditions—sparse rewards, misleading correlations, and input perturbations—where NNs typically struggle.
+5.2.1 Signal Dilution
+Signal dilution occurs when rare events, such as sparse rewards, are statistically overwhelmed by more common non-event data during training (Shyalika et al., 2024; Zhao et al., 2018; He & Cheng, 2021; Tessler et al., 2017).
+In NNs with fixed architectures, shared weights inherently prioritize patterns from dominant classes or frequent signals, causing rare events to be ”averaged out” or under-represented in learned representations.
+This phenomenon is particularly problematic in RL and classification tasks with imbalanced datasets, where the model fails to retain or react to critical but infrequent signals.
+Experimental Setup.
+We compare GBRL with NNs in a linear equation variable isolation environment with sparse rewards, inspired by (Poesia et al., 2021).
+The agent observes the coefficients $(a,b,c)$ of a linear equation $ax+b=c$, and needs to isolate $x$, receiving a large reward only upon completion.
+Actions are multi-discrete, involving the selection of an operation, a digit $\in[0,9]$, and its sign.
+The optimal policy solves the simple task in 2 steps.
+We also evaluate two more complex variations with an additional variable $y$, where the optimal policy requires 3–4 steps.
+For these, we introduce intermediate goals with small rewards to guide the agent.
+Results.
+While simple for humans, these tasks remain challenging for NNs due to signal dilution (Figure 5).
+Although they can make some progress in the simple task ($ax+b=c$), they fail when the equation complexity increases, even when given intermediate rewards.
+In contrast, GBRL converges to and retains a stable (and optimal) policy throughout the process.
+GBTs naturally segment the decision space into explicit rule-based splits, preventing signal dilution while correctly preserving the rare but crucial reward signals.
+Consequently, we see that GBRL can preserve and exploit rare but critical reward signals—evidence that GBT-based methods are inherently less prone to signal dilution.
+Figure 5: Signal dilution, variable isolation task.
+Mean and standard deviation of the average episodic reward during training.
+GBRL was trained for 15M training steps and NN for 30M.
+Episodes are terminated after 50 steps if not solved.
+5.2.2 Spurious Correlations
+Experimental Setup.
+The agent is placed in a grid environment with three balls (red, green, and blue), and a red box.
+The ball locations are randomized.
+The agent observes an objective specifying which ball to pick up.
+We experiment with three box-placement configurations: adjacent to the target ball, adjacent to one of the non-target balls, and without a box in the room.
+This task tests whether the policy overfits spuriously correlated features.
+We then analyze the results using SHAP (Lundberg & Lee, 2017), a method based on Shapley values (ARROW et al., 1953) that measures the contribution of each feature to the final prediction of a model.
+Results.
+GBRL consistently outperforms NNs in all scenarios (Figure 6), demonstrating its ability to ignore irrelevant distractions, such as the box, and reliably reach the target regardless of its placement.
+To understand GBRL’s performance, we analyzed the SHAP values for the agent’s surroundings and the mission feature when predicting the logit for picking up the target object (Figure 7).
+GBRL assigns the strongest positive SHAP value to the target object, followed by the mission string, aligning perfectly with the goal.
+Non-target objects receive the most negative SHAP values, suppressing distractions and reinforcing goal-directed behavior.
+In contrast, NNs assign negative SHAP values to the mission feature and rely heavily on the grid surrounding the target object.
+This indicates a reliance on spurious correlations and a lack of goal-oriented strategies.
+These findings underscore GBRL’s ability to focus on task-relevant features and ignore deceptive cues, reinforcing our hypothesis that GBT’s inductive bias provides robustness against spurious correlations.
+Next, we consider how this robustness extends to other forms of state perturbations.
+Figure 6: Spurious correlations, MiniGrid. Mean and standard deviation of the episodic reward. Plots outlined in green are in-distribution, while plots outlined in red are out-of-distribution.
+(a) GBRL SHAP.
+(b) NN SHAP.
+Figure 7: Spurious correlations revealed by feature importance in MiniGrid.
+The visualization shows normalized SHAP values for the concluding action in a successful evaluation run, revealing the magnitude and direction of each feature’s impact on model predictions.
+In this scenario, a red box was positioned next to the target (red ball), and both models converged to a successful policy.
+GBRL prioritizes the mission, ignores the red box, and contrasts the blue ball.
+In contrast, the NN focuses on the cells surrounding the target, while ignoring the blue ball and contrasting the box and the mission.
+5.2.3 Robustness to State Perturbations
+When encountering an out-of-distribution (OOD) state, a GBT model will always output a result within the same range seen in training.
+In contrast, NNs may extrapolate and exhibit unexpected behavior (Xu et al., 2021; Meinke & Hein, 2020; Ulmer & Cinà, 2021).
+We consider three tasks.
+First, robustness to irrelevant information.
+Second, robustness to random noise.
+And finally, robustness to missing features.
+These experiments aim to test the inherent robustness properties of GBTs when encountering different types of noise during training.
+Irrelevant Information.
+We consider a similar fetch experiment as before (Section 5.2.2).
+Now, a new object is placed in the room at a random location.
+This object is not part of the mission and only interferes with navigation.
+The interfering object differs from the task objects in shape, color, or both.
+Figure 8 illustrates how introducing a distractor impacts policy performance.
+When the environment is free of distracting objects, NNs are able to solve the task, converging to a stable and ideal solution.
+Surprisingly, although the purple-box object shares no features with the task objects (blue, green, and red balls), NNs exhibit a drop in performance but are able to solve the task in a subset of seeds.
+This performance degradation increases when the confusing object shares more features with the goal objects, such as purple-ball and red-box.
+In contrast, GBRL successfully and efficiently solves the task, regardless of the additional irrelevant object placed in the scene.
+Figure 8: Robustness to irrelevant objects, MiniGrid. Mean and standard deviation of the average episodic reward during training. GBRL was trained for 5M training steps, and NN for 20M.
+Random Noise.
+In the classic-control environments (Section 5.1), we train agents in a noiseless setting.
+Then, to evaluate the robustness of the resulting policies, we add random zero-centered Gaussian noise to the input states as defined by: $\operatorname{s}_{t}=\operatorname{s}_{t}+\epsilon|\operatorname{s}_{t}|$.
+The results show that these characteristics translate to practical outcomes (Figure 9).
+In classic-control environments, a trained GBT-based policy is much more robust to small variations in the state space compared to NNs.
+We attribute this performance to how GBTs partition the input space.
+This partitioning occurs on hard thresholds that are split according to single features.
+Although small perturbations may be sufficient to completely change decision outcomes in NNs (Lütjens et al., 2020; Carlini & Wagner, 2017), GBRL exhibits robustness to random additive noise.
+Figure 9: Robustness to input perturbations, classic-control.
+Models were trained without perturbations and then evaluated on noisy observations.
+At each timestep, the observation was perturbed following $\operatorname{s}_{t}=\operatorname{s}_{t}+\epsilon|\operatorname{s}_{t}|$, where $\epsilon\sim\mathcal{N}(0,\text{var})$.
+We report the mean and standard deviation of episodic reward across 100 evaluation episodes for each $\text{Variance}\in[0,0.2]$.
+Missing Features.
+We modify scenarios in the football domain (Section 5.1) to support the random drop of a player from the opposing team.
+Policies are trained on both the original task and a modified variant.
+We then evaluate model robustness by testing how a policy trained against $N$ random players performs when faced with $N+1$, or how a policy trained against a full team behaves when a random opponent is removed.
+The results show that GBRL outperforms NNs in OOD tasks (Figure 10), particularly in the 11 vs 11 academy scenario.
+Notably, GBRL maintains performance when evaluated against 10 opponents and can solve 11 vs.
+11 even when trained on 10.
+In contrast, NNs struggle in all OOD cases.
+This suggests GBRL better isolates features that are essential to solving the task.
+Figure 10: Out-of-distribution evaluation, football.
+Mean and standard deviation of episodic reward across 100 evaluation episodes.
+Plots outlined in green are in-distribution, while plots outlined in red are out-of-distribution.
+6 Ablation Study
+In this section, we examine the sensitivity of GBRL to key hyperparameters.
+We systematically changed one hyperparameter at a time while keeping the others fixed.
+For each experiment, we trained an agent in the CartPole-v1 environment for 1M timesteps using GBT-based PPO.
+Learning Rate Variation Across Policy and Value Function Components.
+We investigated the impact of the learning rate for both the policy and value function components.
+In these experiments, we held one learning rate constant while varying the other.
+Consistent with observations in NN training, we found that excessively high learning rates destabilized the learning process, whereas overly low rates led to significantly slower convergence (as illustrated in Figure 11).
+Figure 11: Ablation: Learning rate variation across policy (left) and value function (right) components.
+Mean and standard deviation of the average episodic reward during training.
+Each subplot shows the effect of varying the learning rate for one component (policy or value function) while keeping the other fixed.
+Limitations of Tree Depth.
+We varied the maximum depth of the decision trees built at each training iteration and evaluated their impact on learning dynamics.
+Figure 12 demonstrates that deeper trees yield more accurate gradient approximations, leading to faster convergence, but also incur a higher computational cost per training step.
+In the paper, we selected a maximum tree depth of 4, balancing convergence speed and wall-clock time.
+Figure 12: Ablation: Effect of tree depth on convergence.
+We varied the maximum depth of the decision trees used in the gradient-boosted models and evaluated their impact on learning dynamics.
+The plot shows the mean and standard deviation of the average episodic reward over time and the total runtime.
+Batch Size Impact on Convergence.
+We analyzed the impact of the batch size on performance.
+In Figure 13, we observe that batch size significantly impacts convergence.
+Specifically, smaller batches result in GBRL building more trees per rollout, improving adaptability.
+However, smaller batches also lead to noisier gradient estimates as a result of limited samples per constructed tree.
+Conversely, larger batches stabilize training by reducing variance through averaging within leaves, as more samples are utilized to construct each tree but build fewer trees per rollout.
+Hence, both excessively small and large batch sizes negatively impact performance.
+Figure 13: Ablation: Batch size impact on convergence.
+Mean and standard deviation of the average episodic reward during training.
+We fix the rollout length to 2048 and vary the batch size.
+Smaller batches improve early learning but suffer from instability, while larger batches stabilize training at the cost of slower convergence.
+Overall, these ablations highlight that GBRL’s performance is sensitive to training frequency, tree depth, and batch size—each of which governs the trade-off between convergence speed and stability.
+While this section focused on general hyperparameters, we further investigate how these trade-offs manifest across different RL algorithms in Appendix B.
+7 Limitations and Future Directions
+GBRL exhibits competitive performance and surprising robustness capabilities.
+However, despite these benefits, we highlight two important limitations.
+(i) Continuous generation of trees.
+As the policy improves through numerous updates, the size of the ensemble increases.
+This unbounded growth has implications for memory usage, computational efficiency, and the feasibility of online real-time adaptation.
+Future work may leverage tree redundancy to produce a more efficient and compact policy.
+For example, strategies for tree pruning, ensemble compression, or dynamically managing ensemble size.
+(ii) Off-policy continuous control.
+Methods like DDPG (Lillicrap et al., 2019) and SAC (Haarnoja et al., 2018) update the policy by differentiating the Q-estimator.
+As GBTs are not differentiable, new solutions are needed to incorporate them into these algorithms.
 8 Conclusion
 Historically, RL practitioners have relied on tabular, linear, and NN-based function approximators.
-But, GBT, a widely successful tool in supervised learning, has been absent from this toolbox.
-We present a method for effectively integrating it into RL and demonstrate domains where it excels compared to NNs.
-GBRL is a step toward solutions that are more interpretable, well suited for real-world tasks with structured data, or capable of deployment on low-compute devices.
-The choice of an RL method depends on the task characteristics: tabular and linear approaches are suitable for small state spaces or simple mappings, while NNs handle complex relationships in unstructured data.
-GBT thrives in complex, yet structured environments.
-In such cases, we observe the advantage of GBRL over NNs, reflecting its already known benefits in supervised learning.
-A crucial component of GBRL is our efficient adaptation of GBT for AC methods, which allows the simultaneous optimization of distinct objectives.
+While GBT evolved as a widely successful tool in supervised learning, it has been absent from this toolbox.
+In this work, we introduced a method for effectively integrating GBT into RL.
+We demonstrated domains where GBT excels compared to NNs, in addition to analyzing its various inherent characteristics.
+We observe that the ideal method depends on the task characteristics.
+Tabular and linear approaches are suitable for small state spaces or simple mappings, while NNs excel when tackling complex relationships in unstructured data.
+In comparison, GBT thrives in complex, yet structured environments.
+In such cases, GBRL’s advantage is highlighted, reflecting GBTs known benefits from supervised learning.
+A crucial component of GBRL is our efficient adaptation of GBT for Actor-Critic methods, which allows the simultaneous optimization of distinct objectives.
 We optimized this approach for large-scale ensembles using GPU acceleration (CUDA).
 Furthermore, GBRL integrates seamlessly with existing RL libraries, promoting ease of use and adoption.
-7 Limitations and Future Directions
-In this work, we integrated the highly popular GBT, typically used in supervised learning, into RL.
-Our results show that GBT is competitive across a range of problems.
-However, we identified several limitations and compelling areas for further research.
-First, a significant challenge lies in the continuous generation of trees.
-As the policy improves through numerous updates, the size of the ensemble increases.
-This unbounded growth has implications for memory usage, computational efficiency, and the feasibility of online real-time adaptation.
-The problem is exacerbated by off-policy methods that build many trees per sample.
-Moreover, the redundancy of trees, especially those from early stages, suggests that the final policy could be represented with a much smaller ensemble.
-Consequently, developing strategies for tree pruning, ensemble compression, or dynamically managing ensemble size could offer crucial optimizations without compromising performance.
-Another key challenge lies in effectively integrating GBT with additional state-of-the-art RL algorithms such as DDPG [23] or SAC [17].
-These require differentiable Q-functions to update the policy.
-Since GBTs are not differentiable, new solutions are needed to incorporate them into these algorithms.
-One such possible direction can be probabilistic trees, where each node represents the probability of traversing the graph.
+GBRL is a step toward well-suited solutions for real-world tasks with structured data.
+Impact Statement
+This paper presents work whose goal is to advance the field of Machine Learning.
+There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
 Appendix
 This appendix provides supplementary materials that support the findings and methodologies discussed in the main text.
-It is organized into four sections to present the full experiment results, implementation details, hyperparameters used during the experiments, training progression plots, and experimental plots, respectively.
+It is organized into five sections to present the full experiment results, implementation details, hyperparameters used during the experiments, training progression plots, and experimental plots, respectively.
 These materials offer detailed insights into the research process and outcomes, facilitating a deeper understanding and replication of the study.
-Appendix A Implementaion Details and Hyperparameters
-Included in this section are implementation details, information regarding compute resources, and tables containing the hyperparameters used in our experiments enabling the reproducibility of our results.
+Appendix A Multi-Objective Training With GBT
+In GBRL, we compute gradients for both the actor and critic objectives per timestep.
+We then concatenate these gradients: $g_{t}:=[g_{t,\text{actor}},g_{t,\text{critic}}]$ and use them to build a decision tree.
+At each candidate node split, we evaluate a score function to decide the best split.
+However, since the actor and critic gradients can differ in magnitude, this score can become biased toward one objective.
+To address this, we explored two strategies:
+  • Gradient normalization per output dimension with L2-based split scoring.
+  • Cosine similarity, which emphasizes gradient direction over magnitude.
+Both approaches are supported in our codebase and produce good results, while the cosine similarity performed slightly better (Figure 14).
+Figure 14: Multi-objective gradient split strategies in PPO-GBRL on CartPole-v1.
+We compare gradient normalization and cosine similarity as methods for handling actor-critic gradient imbalance when building trees.
+Both strategies yield good performance, with cosine similarity achieving slightly faster convergence.
+Curves show average episodic reward over 1M timesteps.
+Appendix B Why Do Certain Algorithms Perform Differently Than Others?
+Apart from inherent algorithmic differences, the main difference for GBRL is the number of fitted trees per update.
+This is analogous to the number of update steps in a neural network, since each tree represents a functional gradient step for the actor or critic.
+To illustrate this effect, we trained agents in the CartPole-v1 environment for 1M timesteps and varied the number of epochs while keeping the batch size and the rollout length fixed, thereby varying the number of updates per rollout.
+We repeated this experiment for several batch sizes while keeping the rollout length fixed at 2048.
+As shown in Figure 15, increasing the number of trees added per rollout improves performance, but increases computational cost.
+Moreover, a larger batch size also increases computational time, due to a larger number of samples used to build each tree.
+Hence, there is a clear trade-off between computational cost and performance.
+This trade-off also explains the difference in performance of various algorithms when combined with GBRL, as their update schemes inherently affect how many trees are fitted per rollout.
+  • A2C - performs a single gradient step per rollout using the entire batch.
+In GBRL, this results in a single constructed tree per rollout.
+Since GBTs update the value and policy functions directly, such large, coarse updates can destabilize learning.
+  • PPO - in contrast, uses minibatches and multiple epochs per rollout to better control the trade-off.
+This allows GBRL to fit several small trees per rollout, yielding more stable and incremental updates.
+  • AWR - is an off-policy algorithm.
+Therefore, GBRL could potentially build many trees without adding sample complexity.
+However, we noticed that our best performing models had a huge computational cost to finish training.
+As a result, we resorted to reducing the number of gradient steps per update at the expense of performance.
+Future work could investigate methods for optimizing GBRL with a fixed tree budget (such as pruning and distillation), which may enable the applicability of more demanding RL algorithms, such as AWR, using GBRL.
+In general, as each tree approximates a gradient step, GBRL benefits from frequent updates to refine its ensemble gradually.
+This makes it particularly well-suited to algorithms that support minibatching and multiple epochs per rollout.
+Figure 15: Effect of number of gradient updates per rollout on PPO-GBRL’s performance in CartPole-v1.
+We fix the batch size and the number of samples collected per rollout, while varying the number of PPO epochs, thereby changing the number of gradient updates per rollout.
+This improves sample efficiency, with more epochs leading to faster convergence.
+Results are shown over 1M timesteps.
+The number of samples collected per rollout was fixed to 2048 for all experiments.
+Appendix C Implementation Details and Hyperparameters
+This section provides implementation details, information regarding compute resources, and tables containing the hyperparameters used in our experiments enabling the reproducibility of our results.
 Table 2 lists GBRL hyperparameters for all experiments.
-A.1 Environments
-The Football domain consists of a vectorized 115-dimensional observation space that summarizes the main aspects of the game and 19 discrete actions.
+C.1 Algorithm Details
+Our GBRL method is implemented within the Stable Baselines3 framework (Raffin et al., 2021), a widely used library for standardized reinforcement learning algorithms.
+For comparisons, we use baseline NNs from RL Zoo3 (Raffin, 2020), which provides tuned hyperparameters per environment.
+For all experiments, we used a standard MLP with two hidden layers, each containing 64 units and tanh activations.
+The hidden layers are shared between the actor and the critic, and each component has an additional separate linear layer.
+This architecture is the default in Stable Baselines3 for the environments we tested and is widely used as a baseline in RL literature.
+We maintained this consistent architecture across all environments to ensure fair comparison.
+A2C (Mnih et al., 2016) is a synchronous, on-policy Actor-Critic algorithm designed to improve learning stability.
+The critic learns a value function, $V(\operatorname{s})$, used to estimate the advantage.
+This advantage is incorporated into the policy gradient updates, reducing variance and leading to smoother learning.
+The policy is updated using the following gradient: $\nabla_{\theta}J(\pi_{\theta})=\mathbb{E}[\nabla_{\theta}\log\pi_{\theta}(\operatorname{a}|\operatorname{s})A(\operatorname{s},\operatorname{a})]$.
+AWR (Peng et al., 2019) is an off-policy Actor-Critic algorithm.
+Provided a dataset $\mathcal{D}$, AWR updates both the policy and the value through supervised learning.
+This dataset can be pre-defined and fixed (offline), or continually updated using the agent’s experience (replay buffer).
+At each training iteration $k$, AWR solves the following two regression problems:
+$V_{k}$ $=\operatorname*{arg\,min}_{V}\mathbb{E}_{\operatorname{s},\operatorname{a}\sim\mathcal{D}}[\|G(\operatorname{s},\operatorname{a})-V(\operatorname{s})\|_{2}^{2}]\,,\,$
+$\pi_{k+1}$ $=\operatorname*{arg\,max}_{\pi}\mathbb{E}_{\operatorname{s},\operatorname{a}\sim\mathcal{D}}[\log\pi(\operatorname{a}|\operatorname{s})\exp(\frac{1}{\beta}A_{k}(\operatorname{s},\operatorname{a}))]\,,$
+where $G(\operatorname{s},\operatorname{a})$ represents the monte-carlo estimate or TD($\lambda$) of the expected return (Sutton & Barto, 2018).
+Advantage computation: We use Generalized Advantage Estimation (GAE) as proposed by Schulman et al.
+(2016) to compute the advantage: $A_{t}=\sum_{l=0}^{\infty}(\gamma\lambda)^{l}\delta_{t+l}$ , where $\delta_{t}=r_{t}+\gamma V(\operatorname{s}_{t+1})-V(\operatorname{s}_{t})$.
+The GAE parameters and for each environment are provided in Table 2 of the appendix.
+Critic: We use on-policy actor-critic algorithms, where the critic predicts the value function (not action-value Q).
+We use a standard L2 loss between the critic’s prediction and the target return estimate: $L_{\text{critic}}(\theta)=||V_{\theta}(\operatorname{s}_{t})-G_{t}||^{2}$, where $G_{t}$ is either a Monte Carlo return estimate or a bootstrapped n-step return, depending on the specific algorithm (PPO, A2C, or AWR).
+C.2 Environments
+Classic-Control & Box2D.
+We evaluate GBRL against NNs on all available classic-control tasks provided in Gymnasium (Towers et al., 2024) and on LunarLander-v2.
+This includes three discrete action space environments: Acrobot-v1, CartPole-v1, MountainCar-v0, and two continuous action space environments: Pendulum-v1 and MountainCarContinuous-v0.
+We trained on both environments for 1M steps (1.5M for LunarLander-v2).
+For NNs we used the hyperparameters provided in RL Baselines3 Zoo (Raffin, 2020), except for AWR where we followed the hyperparameters described in the paper.
+Google Research Football & Atari ram.
+The Football domain (Kurach et al., 2020) consists of a vectorized 115-dimensional observation space that summarizes the main aspects of the game and 19 discrete actions.
 We focus on its academy scenarios, which present situational tasks involving scoring a single goal.
-A standard reward of +1 is granted for scoring, and we employed the "Checkpoints" shaped reward structure.
+A standard reward of +1 is granted for scoring, and we employed the ”Checkpoints” shaped reward structure.
 This structure provides additional points as the agent moves closer towards the goal, with a maximum reward of 2 per scenario.
-The Atari-ram environment consists of a vectorized 128-dimensional observational space representing the 128 byte RAM state and up to 18 discrete actions.
+The Atari-ram domain consists of a vectorized 128-dimensional observational space representing the 128-byte RAM state and up to 18 discrete actions.
 We trained agents in both domains for 10M timesteps.
-The MiniGrid environment [8] is a 2D grid world with goal-oriented tasks requiring object interaction.
+MiniGrid.
+The MiniGrid environment (Chevalier-Boisvert et al., 2023) is a 2D grid world with goal-oriented tasks requiring object interaction.
 The observation space consists of a 7x7 image representing the grid, a mission string, and the agent’s direction.
@@
 All MiniGrid tasks emit a reward of +1 for successful completion and 0 otherwise.
-We trained our NN-based agents on a flattened observation space using the built-in one-hot wrapper.
+We trained our NN-based agents on a flattened observation space using the built-in one-hot wrapper, as specified by RL Baselines3 Zoo.
 For GBRL agents, we generated a 51-dimensional categorical observational space by encoding each unique tile tuple as a categorical string to represent the observed image.
@@
 All agents were trained for 1M timesteps, except for PutNear, FourRooms, and Fetch tasks, which were trained for 10M based on the reported values for PPO NN in RL Baselines3 Zoo.
+Variable Isolation Environment.
+The goal of the variable isolation task is to separate a target variable within a linear equation.
+In the basic task variation, the initial state is a linear equation of the form $ax+b=c$, where $a,b,c$ are digits within the range $[1-9]$ that can be either positive or negative.
+The action space is multi-discrete, where the agent chooses an action type, the digit to use, and the sign of the digit.
+The action types are: addition, subtraction, multiplication, and division.
+The agent receives a reward proportional to the number of steps taken in the episode if it successfully isolates $x$, otherwise the reward is 00.
+In the two-variable linear equation variation, the initial state is a linear equation of the form $ax+by+c=d$, and the goal is to isolate $y$ on the left side.
+The actions are equivalent to the basic task, with the addition or subtraction of a digit multiplied by $x$.
+Finally, in the balanced two-variable linear equation variation, the initial state is a linear equation of the form $ax+b=cy+d$, and the goal is to isolate $y$ on the left side.
+The actions are equivalent to the two-variable linear equation variation with the addition of the operations involving $y$.
+In all the environment variations, the agent observes the coefficients as inputs.
+For the two variable variations, we add a small reward bonus if the agent successfully moves a variable or a constant to the correct side.
+We trained GBRL for 15M steps and NNs for 30M steps.
 batch size clip range ent coef gae lambda gamma
@@
 MiniGrid 512 0.2 0.0 0.95 0.99 20 256 16 0.17 0.01
+Variable Isolation 128 0.62 0.08 0.95 0.99 20 256 16 0.18 0.006
 (a) PPO.
-For continuous action spaces we used log std init = -2 and log std lr = lin_0.0017.
+For continuous action spaces, we used log std init = -2 and log std lr = lin_0.0017.
 We utilized gradient norm clipping for Gym environments.
@@
 Table 2: GBRL hyperparameters - NN represented by an MLP with two hidden layers.
-A.2 Compute Resources
+C.3 Compute Resources
 All experiments were done on the NVIDIA NGC platform on a single NVIDIA V100-32GB GPU per experiment.
@@
 Additionally, the research project involved preliminary experiments and hyperparameter tuning, which required an estimated additional 168 GPU hours.
-Appendix B Detailed Results Tables
+Appendix D Detailed Result Tables
 This section contains tables presenting the mean and standard deviation of the average episode reward for the final 100 episodes within each experiment.
-More specifically, Table 2 presents results for Continuous Control & Block2D environments, Tables 3 and 4 present results for the high-dimensional vectorized environments, and Table 5 presents results for the categorical environments.
-Table 2: Continuous-Control and Box2D environments: Average episode reward for the final 100 episodes.
+More specifically, Table 3 presents results for Continuous Control & Block2D environments, Tables 4 and 5 present results for the high-dimensional vectorized environments, and Table 6 presents results for the categorical environments.
+Table 3: Continuous-Control and Box2D environments: Reported values are the mean and standard deviation of the average episode reward over the final 100 training episodes, aggregated across 5 random seeds.
 Acrobot CartPole LunarLander MountainCar MountainCar Continuous Pendulum-v1
@@
 GBRL: PPO $-87.82\pm 2.16$ $500.00\pm 0.0$ $248.72\pm 59.10$ $-110.55\pm 15.60$ $89.42\pm 5.73$ $-246.89\pm 20.61$
-Table 3: Football Academy environments: Average episode reward for the final 100 episodes.
+Table 4: Football Academy environments: Reported values are the mean and standard deviation of the average episode reward over the final 100 training episodes, aggregated across 5 random seeds.
 3 vs 1 with keeper Corner Counterattack Easy Counterattack Hard Empty Goal Empty Goal Close
@@
 GBRL: PPO $1.87\pm 0.09$ $1.85\pm 0.08$ $1.83\pm 0.04$ $1.95\pm 0.02$ $1.73\pm 0.06$
-Table 4: Atari-ramNoFrameskip-v4 environments: Average episode reward for the final 100 episodes.
+Table 5: Atari-ramNoFrameskip-v4 environments: Reported values are the mean and standard deviation of the average episode reward over the final 100 training episodes, aggregated across 5 random seeds.
 Alien Amidar Asteroids Breakout Gopher
@@
 GBRL: PPO $2160.8\pm 826.92$ $6888.66\pm 756.18$ $2069.22\pm 538.62$ $15.40\pm 6.55$ $434.84\pm 31.83$
-Table 5: MiniGrid environments: Average episode reward for the final 100 episodes.
+Table 6: MiniGrid environments: Reported values are the mean and standard deviation of the average episode reward over the final 100 training episodes, aggregated across 5 random seeds.
 DoorKey-5x5 Empty-Random-5x5 Fetch-5x5-N2 FourRooms GoToDoor-5x5
@@
 GBRL: PPO $0.95\pm 0.00$ $0.44\pm 0.19$ $0.88\pm 0.02$ $0.97\pm 0.00$
-Appendix C Training Plots
+Appendix E Learning Curves
 This section presents learning curves depicting model performance throughout the training phase.
-Figures 8, 9, 10 and 11 show the training reward as a function of environment steps of the agents trained in the experiments.
+Figures 16, 17, 18 and 19 show the training reward as a function of environment steps of the agents trained in the experiments.
 The column order is: A2C, AWR, and PPO.
 Figure 16: Classic Control and Box2D environments: Training reward as a function of environment steps.
+Reported values are the mean and standard deviation of the average episode reward over the final 100 training episodes, aggregated across 5 random seeds.
 Figure 17: Football Academy environments: Training reward as a function of environment step.
+Reported values are the mean and standard deviation of the average episode reward over the final 100 training episodes, aggregated across 5 random seeds.
 Figure 18: Atari-ramNoFrameskip-v4 environments: Training reward as a function of environment step.
+Reported values are the mean and standard deviation of the average episode reward over the final 100 training episodes, aggregated across 5 random seeds.
 Figure 19: MiniGrid environments: Training reward as a function of environment step.
+Reported values are the mean and standard deviation of the average episode reward over the final 100 training episodes, aggregated across 5 random seeds.
 Figure 20: Sharing actor critic tree structure significantly increases efficiency while retraining similar performance.